### v2

# Approaches to critical point theory via sequential and parametrized topological complexity

Abstract.
The Lusternik-Schnirelmann category of a space was introduced to obtain a lower bound on the number of critical points of a $C^{1}$-function on a given manifold. Related to Lusternik-Schnirelmann category and motivated by topological robotics, the topological complexity (TC) of a space is a numerical homotopy invariant whose topological properties are an active field of research. The notions of sequential and parametrized topological complexity extend the ideas of topological complexity. While the definition of TC is closely related to Lusternik-Schnirelmann category, the connections of sequential and parametrized TC to critical point theory have not been fully explored yet. In this article we apply methods from Lusternik-Schnirelmann theory to establish various lower bounds on numbers of critical points of functions in terms of sequential and parametrized TCs. We carry out several consequences and applications of these bounds, among them a computation of the parametrized TC of the unit tangent bundles of $(4m-1)$-spheres.
Introduction
Relating topological invariants of spaces to numbers of critical points of functions is one of the central ideas in topology and geometry. In the 1920s, Lusternik and Schnirelmann introduced their notion of category of a space to find a lower bound on these numbers for arbitrary $C^{1}$-functions on a closed smooth manifold. Since then, there have been many extensions and applications of Lusternik-Schnirelmann category to critical point theory, e.g. to infinite-dimensional Banach manifolds by R. Palais in [Pal66b] and to equivariant settings by T. Bartsch in [Bar93] as well as M. Clapp and D. Puppe in [CP91].
Another common approach to relate topological properties of manifolds to critical points of functions is Morse theory, see [Mil63]. While Morse theory often produces stronger bounds on the numbers of critical points, its downside is that the function under investigation needs to be of class $C^{2}$ and its critical points need to fulfill a non-degeneracy condition, which is only satisfied generically.
In recent years, invariants similar to Lusternik-Schnirelmann category have been studied in topological robotics. Motivated by the motion planning problem from robotics, M. Farber introduced the notion of topological complexity of a topological space $X$ in [Far03] and developed it further in several subsequent articles. For a path-connected topological space $X$, let $PX=C^{0}([0,1],X)$ be the path space endowed with the compact-open topology and let
$\pi\colon PX\to X\times X,\quad\pi(\gamma)=(\gamma(0),\gamma(1)),$
be the path fibration. The topological complexity of $X$ is defined as the smallest integer $k\in\mathbb{N}$ such that there exists an open cover $U_{1},U_{2},\ldots,U_{k}$ of $X\times X$ with $U_{i}$ admitting a continuous section of $\pi$ for each $i\in\{1,2,\ldots,k\}$. Here, we note that we use Farber’s original definition instead of the reduced definition which is by one smaller than ours and which has become very common in the literature.
Since M. Farber’s seminal work, many results on and many variants of topological complexity have been studied. In this article, we shall consider the sequential or higher topological complexity of a space $X$, denoted by $\operatorname{\mathsf{TC}}_{r}(X)$ and the (sequential) parametrized topological complexity of a fibration $p\colon E\to B$, denoted by $\operatorname{\mathsf{TC}}_{r}[p\colon E\to B]$.
We want to briefly discuss these two concepts. Let $r\in\mathbb{N}$ with $r\geq 2$ and let $X$ be a path-connected space. Then
$\pi_{r}\colon PX\to X^{r},\quad\pi_{r}(\gamma)=\big{(}\gamma(0),\gamma(\tfrac{1}{r-1}),\ldots,\gamma(\tfrac{r-2}{r-1}),\gamma(1)\big{)},$
is a fibration. The $r$-th sequential topological complexity of $X$ is defined as the smallest integer $k\in\mathbb{N}$ such that there exists an open cover $U_{1},U_{2},\ldots,U_{k}$ of $X^{r}$ such that each $U_{i}$, $i\in\{1,2,\ldots,k\}$ admits a continuous section of $\pi_{r}$. This notion was defined by Y. Rudyak in [Rud10]. Note that for $r=2$, we obtain that $\operatorname{\mathsf{TC}}_{2}(X)=\operatorname{\mathsf{TC}}(X)$, the topological complexity of $X$.
The notion of parametrized topological complexity was introduced by D. Cohen, M. Farber and S. Weinberger in [CFW21] and generalized to sequential parametrized topological complexities by M. Farber and A. Paul in [FP22]. It has since been further studied by the same authors in [FP23] and by M. Farber and J. Oprea in [FO24].
Given $r\in\mathbb{N}$ with $r\geq 2$ and a fibration $p\colon E\to X$ with path-connected fibers, we consider
$E_{X}^{r}=\{(u_{1},u_{2},\ldots,u_{r})\in E^{r}\mid p(u_{1})=p(u_{2})=\ldots=p(u_{r})\}$
and
$E_{X}^{I}=\{\gamma\in PE\mid p(\gamma(t))=p(\gamma(0))\,\,\,\text{for all}\,\,\,t\in[0,1]\}.$
As a consequence of the main result of [CFW22, Appendix], the map
$\Pi_{r}\colon E_{X}^{I}\to E_{X}^{r},\qquad\Pi_{r}(\gamma)=\big{(}\gamma(0),\gamma(\tfrac{1}{r-1}),\ldots,\gamma(\tfrac{r-2}{r-1}),\gamma(1)\big{)},$
is a fibration. The $r$-th sequential parametrized topological complexity of $p\colon E\to X$, denoted by $\operatorname{\mathsf{TC}}_{r}[p\colon E\to X]$ is then defined as the smallest value of $k\in\mathbb{N}$ such that there exists an open cover $U_{1},U_{2},\ldots,U_{k}$ of $E_{X}^{r}$ with each $U_{i}$, $i\in\{1,2,\ldots,k\}$ admitting a continuous section of $\Pi_{r}$.
The definitions of topological complexity and its variants are strongly related to Lusternik-Schnirelmann category, in fact all of these concepts are special cases of the notion of sectional category of a fibration, which was introduced by A. Schwarz in [Sch66] under the name of genus. However, in contrast to Lusternik-Schnirelmann category, the relations between topological complexity and its variants on the one hand and critical point theory on the other hand has only been studied in a few instances. In [Far08, Definition 4.30], M. Farber defined the notion of a navigation function on a closed manifold $M$ to establish upper bounds on $\operatorname{\mathsf{TC}}(M)$ in terms of critical points of functions on $M\times M$. There, a function $F\colon M\times M\to\mathbb{R}$ is called a navigation function if it is non-negative, if $F^{-1}(\{0\})=\Delta_{M}$, where $\Delta_{M}\subset M\times M$ is the diagonal, and if it is a Morse-Bott function. By [Far08, Theorem 4.32], it holds for such a navigation function that
$\operatorname{\mathsf{TC}}(M)\leq\sum_{i=1}^{k}\max\{\operatorname{\mathsf{TC}}_{M}(C)\ |\ C\subset M\times M\,\,\,\text{is a critical manifold with}\,\,\,F(C)=c_{i}\}+1,$
where $c_{1},\dots,c_{k}$ are the non-zero critical values of $F$. Here, $\operatorname{\mathsf{TC}}_{M}(C)$ is the subspace complexity of the set $C\subset M\times M$ as defined in [Far08, Definition 4.20], a term that we shall discuss below in greater generality. Navigation functions on lens spaces were investigated by A. Costa in [Cos10].
In this article we generalize the aforementioned approach to topological complexity via critical point theory in several ways.
  • We follow the topological approach to Lusternik-Schnirelmann theory of Y. Rudyak and F. Schlenk from [RS03] to establish results on numbers of fixed points of self-maps homotopic to the identity. This generalizes the situation for critical points, which corresponds to the case of a time-1 map of a negative gradient flow.
  • We study upper bounds on sequential and parametrized topological complexities of sublevel sets in terms of critical points for arbitrary $C^{1}$-functions and obtain results for generalized navigation functions as special cases.
  • In contrast to Farber’s navigation functions, we do not make any assumptions of non-degeneracy on the critical points of the functions.
Throughout the following, given a function $F\colon X\to\mathbb{R}$, we let $F^{\lambda}:=F^{-1}((-\infty,\lambda])$ denote the corresponding closed sublevel set for each $\lambda\in\mathbb{R}$. The following theorem is our most general result for sequential topological complexities. All of the terminology used in these assertions will be made precise below. For the moment, we just note that condition (D) mentioned in these statements is a topological generalization of the Palais-Smale condition.
Theorem (Theorem 2.10).
Let $r\in\mathbb{N}$ with $r\geq 2$ and let $X$ be a path-connected ANR. Let $\varphi\colon X^{r}\to X^{r}$ be continuous and homotopic to the identity, let $F\colon X^{r}\to\mathbb{R}$ be a Lyapunov function for $\varphi$ which is bounded from below. Assume that $F(\mathrm{Fix\;}\varphi)$ is discrete and that $(\varphi,F)$ satisfies condition (D).
  1. a)
    For all $\lambda\in\mathbb{R}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,X}(F_{\mu}\cap\mathrm{Fix\;}\varphi).$
  2. b)
    If $\lambda\in\mathbb{R}$ is such that $\mathcal{C}_{\mu}(\varphi,F)$ is finite for all $\mu\leq\lambda$, then
    $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(\varphi,F)\}.$
We further establish a similar result for sequential parametrized topological complexity, where we need to put the additional assumption that the self-map is fiber-preserving for a certain fibration.
Theorem (Theorem 5.12).
Let $r\in\mathbb{N}$ with $r\geq 2$, let $E$ and $X$ be ANRs and let $p\colon E\to X$ be a fibration with path-connected fibers. We denote $E^{r}_{X}=\{(u_{1},\dots,u_{r})\in E^{r}\mid p(u_{1})=\dots=p(u_{r})\}$ and consider
$q_{r}\colon E^{r}_{X}\to X,\quad q_{r}(u_{1},\dots,u_{r})=p(u_{1}).$
Let $\Phi\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ be continuous with $\Phi(u,0)=u$ for all $u\in E^{r}_{X}$ and such that
$q_{r}(\Phi(u,t))=q_{r}(u)\quad\forall u\in E^{r}_{X},\ t\in[0,1].$
Let $F\colon E^{r}_{X}\to\mathbb{R}$ be a Lyapunov function for $\Phi_{1}:=\Phi(\cdot,1)$ which is bounded from below and such that $(\Phi_{1},F)$ satisfies condition (D). Then
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\mathrm{Fix\;}\Phi_{1})\qquad\forall\lambda\in\mathbb{R}.$
As explained above, these general results include the case of critical points of $C^{1}$-functions. Applying the two theorems to time-1 maps of negative gradient flows, we obtain the following consequences.
Theorem (Corollary 2.11).
Let $(X,g)$ be a connected and complete Riemannian Banach manifold and let $F\in C^{1}(X^{r})$ be bounded from below. Assume that $(F,g^{r})$ satisfies the Palais-Smale condition and that $F(\operatorname{Crit}F)\subset\mathbb{R}$ is discrete. For all $\lambda\in\mathbb{R}$ it holds that
$\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}.$
For the setting of sequential parametrized topological complexity the arguments are more involved, since the gradient flow of an arbitrary function $F\colon E_{X}^{r}\to\mathbb{R}$ does not need to respect the fibers of the fibration $q_{r}\colon E_{X}^{r}\to X$. To overcome this problem we introduce the notion of a vertically proportional vector field on $E_{X}^{r}$ and show the following.
Theorem (Theorem 6.5).
Let $E$ and $X$ be paracompact Banach manifolds and let $p\colon E\to X$ be a smooth fibration with path-connected fibers and a surjective submersion. Let $F\in C^{1}(E^{r}_{X})$ be bounded from below. Assume that $(F,\bar{g})$ satisfies the Palais-Smale condition for a suitable complete metric $\bar{g}$ on $E^{r}_{X}$ and that $\nabla^{\bar{g}}F$ is vertically proportional with respect to $q_{r}\colon E^{r}_{X}\to X$ and $\bar{g}$. For each $\lambda\in\mathbb{R}$ it holds that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\left\{\operatorname{\mathsf{TC}}_{r,p}(C)\mid C\in\mathcal{C}_{\mu}(F)\right\}.$
We then generalize Farber’s notion of navigation function to the sequential and the sequential parametrized setting and apply the above results to the thus-obtained functions. As a geometric application, we derive the following result from the study of navigation functions in the case of $r=2$.
Theorem (Theorem 3.5).
Let $n\in\mathbb{N}$, let $M\subset\mathbb{R}^{n}$ be a compact hypersurface and let
$\alpha(M):=\#\left\{\{x,y\}\subset M\ \middle|\ x\neq y,\ T_{x}M=T_{y}M=(x-y)^{\perp}\right\}.$
Then
$\alpha(M)\geq\operatorname{\mathsf{TC}}(M)-1.$
In particular, there are at least $\operatorname{\mathsf{TC}}(M)-1$ distinct unordered pairs of points $\{x,y\}\subset M$ which satisfy $T_{x}M=T_{y}M$.
We further consider functions of the form
$F\colon X^{r}\to\mathbb{R},\qquad F(x_{1},\ldots,x_{r})=\sum_{i=1}^{r}f(x_{i}),$
where $f\colon X\to\mathbb{R}$ is a continuously differentiable function on the smooth manifold $X$. We will derive upper bounds for the sequential topological complexities $\operatorname{\mathsf{TC}}_{r}(f^{\lambda})$ of the closed sublevel set $f^{\lambda}=f^{-1}((-\infty,\lambda])$. Note that in contrast to the above results, here we obtain bounds for the sequential topological complexities of $f^{\lambda}$ instead of their subspace versions with respect to the ambient space $X$. An analogous, but technically more intricate, result is derived for sequential parametrized topological complexities. We will then apply the latter to a framework of equivariant fiber bundles in which the assumptions become more tangible.
Theorem (Corollary 8.8).
Let $G$ be a connected Lie group, let $E$ and $B$ be closed $G$-manifolds, such that the $G$-action on $B$ is transitive, and let $p\colon E\to B$ be a $G$-equivariant fiber bundle. Then for any $G$-invariant $f\in C^{1}(E)$ it holds that
$\operatorname{\mathsf{TC}}_{r}[p\colon E\to B]\leq\sum_{\mu\in\mathbb{R}}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{B}\dots\times_{B}\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \mu_{1},\dots,\mu_{r}\in\mathbb{R},\,\sum_{i=1}^{r}\mu_{i}=\mu\Big{\}}.$
As an application of this result, we will obtain the following:
Theorem (Theorem 8.10).
Let $m\in\mathbb{N}$ and let $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ be the unit tangent bundle of the standard $(4m-1)$-sphere $\mathbb{S}^{4m-1}\subset\mathbb{R}^{4m}$. Then for all $r\in\mathbb{N}$ with $r\geq 2$ it holds that
$\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}]=r+1.$
This article is organized as follows. In section 1, we recall the topological approach to Lusternik-Schnirelmann theory from [RS03], in particular the notion of an index function, and its connection to the Palais-Smale condition. We introduce sequential subspace complexities in section 2, establish their index function properties and carry out the general consequences of the Lusternik-Schnirelmann theorem for sequential topological complexities. The consequences for critical points of $C^{1}$-functions are carried out in that section as well. These results are then applied to navigation functions in section 3 and to product functions in section 4, yielding the abovementioned consequences and applications. In section 5 we introduce sequential parametrized subspace complexities and establish a setting in which they form index functions to which the Lusternik-Schnirelmann-type approach can be applied. Section 6 focuses on the notion of vertical proportionality of a vector field. After introducing this condition, we apply the results of section 5 to differentiable functions with vertically proportional gradients and conclude with a discussion of parametrized navigation functions. In section 7, we will study the parametrized TC of restrictions of fibrations to sublevel sets of functions on their total spaces under the assumption that these are again fibrations and by applying the main results of section 6 to another special class of functions. To obtain examples of fibrations to which this approach applies, we study equivariant fiber bundles over homogeneous spaces in section 8 and establish an upper bound for the parametrized TC of equivariant fiber bundles between closed manifolds with homogeneous codomains. Finally, we apply this result to unit tangent bundles of odd-dimensional spheres and compute their parametrized topological complexities in certain cases.
Notation
We will use the following conventions throughout the manuscript. Let $X$ be a topological space and let $f\colon X\to\mathbb{R}$ be a function.
  • Let $\mathfrak{P}(X)$ denote the power set of $X$.
  • Given a map $\varphi\colon X\to X$, we denote its fixed point set by
    $\mathrm{Fix\;}\varphi=\{x\in X\mid\varphi(x)=x\}.$
  • We denote the closed sublevel and level sets of $f$ , resp., by
    $f^{a}:=\{x\in X\mid f(x)\leq a\},\quad f_{a}:=\{x\in X\mid f(x)=a\}\quad\forall a\in\mathbb{R}.$
  • If $X$ is a smooth manifold and $f$ is a differentiable function, then we denote
    $\operatorname{Crit}f=\{x\in X\ |\ Df_{x}=0\},\qquad\operatorname{Crit}_{\mu}f:=\operatorname{Crit}f\cap f_{\mu}\quad\forall\mu\in\mathbb{R}.$
  • Given $r\in\mathbb{N}$ with $r\geq 2$ we put
    $\Delta_{r,X}:=\{(x,x,\dots,x)\in X^{r}\mid x\in X\}.$
  • Given $A\subset X$, we call a map $\Phi\colon A\times[0,1]\to X$ a deformation of $A$ if it is a homotopy from the inclusion $A\hookrightarrow X$ to another map.
  • Given a vector field $Y\colon M\to TM$ on a manifold $M$, we put
    $\mathrm{sing}(Y):=\{x\in M\mid Y(x)=0\}.$
Acknowledgements
M. S. was partially funded by the Deutsche Forschungsgemeinschaft (German Research Foundation) – grant agreement number 518920559. M. S. is grateful for the support by the Danish National Research Foundation through the Copenhagen Centre for Geometry and Topology (DNRF151).
1. Topological Lusternik-Schnirelmann theory
In this section, we want to recall the topological approach to Lusternik-Schnirelmann theory which was carried out by Rudyak and Schlenk in [RS03]. The basic notions required to develop their approach are summarized in the following definition.
Definition 1.1.
Let $X$ be a topological space and let $\varphi\colon X\to X$ be continuous.
  1. a)
    An index function on $X$ is a map $\nu\colon\mathfrak{P}(X)\to\mathbb{N}_{0}\cup\{+\infty\}$ which has the following properties:
    1. (i)
      (Monotonicity)  If $A\subset B\subset X$, then $\nu(A)\leq\nu(B)$.
    2. (ii)
      (Subadditivity)  For all $A,B\subset X$ it holds that
      $\nu(A\cup B)\leq\nu(A)+\nu(B).$
    3. (iii)
      (Continuity)  If $A\subset X$ is closed, then there exists an open neighborhood $U$ of $A$ with $\nu(A)=\nu(U)$.
  2. b)
    An index function $\nu$ on $X$ is called $\varphi$-supervariant if
    $\nu(\varphi(A))\geq\nu(A)$
    for all closed subsets $A\subset X$.
  3. c)
    A continuous function $f\colon X\to\mathbb{R}$ is called a Lyapunov function for $\varphi$ if
    $f(\varphi(x))\leq f(x)\qquad\forall x\in X.$
Remark 1.2.
  1. (1)
    The purpose of the notion of an index function is an axiomatization of the crucial properties of relative Lusternik-Schnirelmann categories. Namely, as shown in [CLOT03, Chapter 1], given a path-connected $X$ we obtain an index function on $X$ by putting
    $\nu(A):=\operatorname{\mathsf{cat}}_{X}(A),$
    i.e. the minimal value of $n\in\mathbb{N}$, such that $A$ can be covered by the union of $n$ open subsets of $X$ whose inclusions are nullhomotopic.
  2. (2)
    Our definition of an index function is actually a slight simplification of the one from [RS03], in which index functions are defined relative to a certain subset. A similar notion is introduced under the name of an abstract category in [CLOT03, Section 1.7].
  3. (3)
    In [RS03], the $\varphi$-supervariance of an index function is defined with respect to arbitrary subsets of $X$, not only closed ones. However, a careful check of the proofs in [RS03, Section 2] shows that it is sufficient to assume supervariance for closed subsets of $X$ to obtain the results that we make use of.
Rudyak and Schlenk have further introduced a topological condition that mimicks the classical Palais-Smale (PS) condition. Its relation to the original PS condition will be clarified below.
Definition 1.3 ([RS03, Definition 1.1]).
Let $X$ be a topological space, let $\varphi\colon X\to X$ be continuous and let $f\colon X\to\mathbb{R}$ be a Lyapunov function for $\varphi$. We say that $(\varphi,f)$ satisfies condition (D) if the following two conditions hold:
  1. ($D_{1}$)
    $f$ is a strong Lyapunov function for $\varphi$, i.e. it holds that
    $f(\varphi(x))<f(x)\qquad\forall x\in X\smallsetminus\mathrm{Fix\;}\varphi.$
  2. ($D_{2}$)
    If $A\subset X$ is chosen such that $f$ is bounded on $A$ and such that
    $\inf\{f(a)-f(\varphi(a))\mid a\in A\}=0,$
    then $\overline{A}\cap\mathrm{Fix\;}\varphi\neq\emptyset$.
The topological abstraction of the classical Lusternik-Schnirelmann theorem by Rudyak and Schlenk is the following.
Theorem 1.4 ([RS03, Theorem 2.3]).
Let $X$ be a topological space, let $\varphi\colon X\to X$ be continuous and let $\nu\colon\mathfrak{P}(X)\to\mathbb{N}\cup\{+\infty\}$ be a $\varphi$-supervariant index function. Assume that there exists a Lyapunov function $f\colon X\to\mathbb{R}$, which is bounded from below, such that $(\varphi,f)$ satisfies condition (D) and that $f(\mathrm{Fix\;}\varphi)\subset\mathbb{R}$ is discrete. Then
$\nu(f^{a})\leq\sum_{\lambda\in(-\infty,a]}\nu(f_{\lambda}\cap\mathrm{Fix\;}\varphi)\qquad\forall a\in\mathbb{R}.$
The original version of the Lusternik-Schnirelmann theorem and in most of its generalizations are re-obtained from Theorem 1.4 by letting $X$ be a differentiable manifold and choose $\varphi$ as the time-$1$ map of a negative gradient flow of a $C^{1}$-function $f\colon X\to\mathbb{R}$. Then $\mathrm{Fix\;}\varphi=\operatorname{Crit}f$, so that Theorem 1.4 can be used to estimate the number of critical points of a function in one of its sublevels from below.
Regarding this setting, we want to recall the PS condition and discuss its relation to condition (D). The following definition of a pseudo-gradient is a mild generalization of [Str08, Definition II.3.8].
Definition 1.5.
[…]
  1. (i)
    $X$ is Lipschitz continuous with respect to $g$ and the Sasaki metric on $TM$ induced by $g$, see e.g. [Lan99, Section X.4] for a description of the latter.
  2. (ii)
    There exists $C_{1}>0$, such that
    $\|X(x)\|\leq C_{1}\cdot\min\{\|\nabla^{g}F(x)\|,1\}\qquad\forall x\in M.$
  3. (iii)
    There exists $C_{2}>0$, such that
    $DF_{x}(X(x))\geq C_{2}\cdot\min\{\|\nabla^{g}F(x)\|,1\}\cdot\|\nabla^{g}F(x)\|\qquad\forall x\in M.$
Remark 1.6.
Let $(M,g)$ be a Riemannian Banach manifold.
  1. (1)
    By a fundamental theorem of K. Nomizu and H. Ozeki, see [NO61], a differentiable Banach manifold admits a complete Riemannian metric if and only if it is paracompact. Thus, working with complete Riemannian Banach manifolds implicitly implies that the manifold under consideration is paracompact.
  2. (2)
    It is shown in [Str08, Lemma II.3.9], see also [Pal66b], that any $F\in C^{1}(M)$ admits a pseudo-gradient vector field. Moreover, by basic results on ordinary differential equations on manifolds, if $(M,g)$ is complete, then any pseudo-gradient for $F$ admits a complete flow.
  3. (3)
    Let $F\in C^{1}(M)$ and let $X\colon M\to TM$ be a pseudo-gradient of $F$. It follows from properties (ii) and (iii) of Definition 1.5 that $\mathrm{sing}(X)=\operatorname{Crit}F$.
The authors of [RS03] study a particular construction of a pseudo-gradient vector field that we shall briefly elaborate upon.
Example 1.7.
Let $(M,g)$ be a complete Riemannian Banach manifold, let $F\in C^{1}(M)$ and let $\nabla F$ denote the gradient of $F$ with respect to $g$. Consider a smooth and monotonically increasing function $\rho\colon\mathbb{R}\to\mathbb{R}$ with the following properties:
  • $\rho(x)=1$ if $x\leq 1$,
  • $\rho(x)\leq x$ if $x\in[1,2]$,
  • $\rho(x)=x$ if $x\geq 2$.
Put $h\colon M\to\mathbb{R}$, $h(x):=\frac{1}{\rho(\|\nabla F(x)\|)}$. Then
$X\colon M\to TM,\qquad X(x):=h(x)\cdot\nabla F(x),$
is a pseudo-gradient vector field for $F$: property (i) from Definition 1.5 follows from the boundedness of $X$ that is inherent in its definition. Considering property (ii) we compute for $x\in M$ with $\|\nabla F(x)\|\leq 2$ that $h(x)\leq 1$ and thus
$\|X(x)\|=h(x)\cdot\|\nabla F(x)\|\leq\|\nabla F(x)\|\leq 2\min\{\|\nabla F(x)\|,1\}.$
For $x\in M$ with $\|\nabla F(x)\|\geq 2$, we obtain $h(x)=\frac{1}{\|\nabla F(x)\|}$ and thus
$\|X(x)\|=1\leq 2\min\{\|\nabla F(x)\|,1\}.$
Concerning property (iii), we compute for each $x\in M$ that
$DF_{x}(X(x))=h(x)\cdot DF_{x}(\nabla F(x))=h(x)\cdot\|\nabla F(x)\|^{2}.$
By definition, $h(x)\geq\min\{\frac{1}{\|\nabla^{g}F(x)\|},1\}$ for each $x\in M$, so it follows that
$DF_{x}(X(x))\geq\min\big{\{}\|\nabla F(x)\|,\|\nabla F(x)\|^{2}\big{\}}\qquad\forall x\in M.$
Hence, $X$ is a pseudo-gradient vector field for $F$.
In [RS03] it is shown that if $f\colon M\to\mathbb{R}$ is any function on a complete Riemannian manifold which satisfies the condition (C) of Palais and Smale, see e.g. [Pal66b], then the time-1 map $\phi_{1}$ of the flow of a negative pseudo-gradient given as in Example 1.7 together with $f$ itself satisfies condition (D). Instead of condition (C), we want to use a different condition in this article, which is more commonly used in studies of gradient flows.
Definition 1.8.
[…]
$\lim_{n\to\infty}\nabla^{g}F(x_{n})=0$
has a convergent subsequence.
The PS condition is slightly stricter than condition (C), see e.g. [MW10, Section 3]. We will make occasional use of the following property.
Lemma 1.9.
Let $(M,g)$ be a complete Riemannian manifold and let $f\in C^{1}(M)$, such that $(f,g)$ satisfies the PS condition. Then $f_{\mu}\cap\operatorname{Crit}f$ has finitely many components for each $\mu\in\mathbb{R}$.
Proof.
Assume there exists some $\mu_{0}\in\mathbb{R}$, such that $N:=f_{\mu_{0}}\cap\operatorname{Crit}f$ has infinitely many components. Then there exists a sequence $(x_{n})_{n\in\mathbb{N}}$ in $N$, for which no two elements lie in the same component of $N$. Evidently, $(x_{n})_{n\in\mathbb{N}}$ does not have a convergent subsequence. But by definition of $N$,
$f(x_{n})=\mu_{0},\qquad\nabla^{g}f(x_{n})=0\qquad\forall n\in\mathbb{N},$
so the PS condition yields that it must have a convergent subsequence. This is a contradiction, so that such a $\mu_{0}\in\mathbb{R}$ can not exist. ∎
The following result is a slight variation of [RS03, Proposition 9.1]. There, the authors study pseudo-gradients of the kind constructed in Example 1.7 as well as condition (C) instead of the PS condition. Their arguments can be transferred easily to the setting considered in this article. For the convenience of the reader, we will carry out the proof in detail.
Theorem 1.10.
Let $(M,g)$ be a complete Riemannian Banach manifold, let $F\in C^{1}(M)$ be bounded from below and assume that $(F,g)$ satisfies the PS condition. Let $X\colon M\to TM$ be a pseudo-gradient vector field of $F$ and let $\phi_{1}\colon M\to M$ be the time-1-map of the flow of $-X$. Then $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D).
Proof.
Let $\phi\colon M\times\mathbb{R}\to M$ be the flow of $-X$ which is well-defined by Remark 1.6.(1) and satisfies $\phi(x,1)=\phi_{1}(x)$ for each $x\in M$. We derive from Remark 1.6.(3) that $\mathrm{Fix\;}\phi_{1}=\mathrm{sing}\,X=\operatorname{Crit}F$. Moreover, we compute for each $x\in M$ that
[…]
and derive from property (iii) of Definition 1.5 that
(1.1) $F(x)\geq F(\phi_{1}(x))\qquad\forall x\in M,$
where equality holds if and only if $DF_{\phi(x,t)}(X(\phi(x,t)))=0$ for all $t\in[0,1]$. By property (iii) of a pseudo-gradient, this is satisfied if and only if $x\in\operatorname{Crit}F=\mathrm{Fix\;}\phi_{1}$. Thus, $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies $(D_{1})$.
In the following let $\nabla^{g}F$ denote the gradient of $F$ with respect to $g$. Let $A\subset M$ be non-empty, such that $F$ is bounded on $A$ and such that $\inf\{F(a)-F(\phi_{1}(a))\mid a\in A\}=0$. As in the proof of [RS03, Proposition 9.1], one obtains sequences $(x_{n})_{n\in\mathbb{N}}$ in $A$ and $(t_{n})_{n\in\mathbb{N}}$ in $[0,1]$, such that
(1.2) $DF_{\phi(x_{n},t_{n})}(X(\phi(x_{n},t_{n})))<\frac{1}{n}\qquad\forall n\in\mathbb{N}.$
Put $y_{n}:=\phi(x_{n},t_{n})$ for each $n$. We derive from (1.2) and property (iii) of a pseudo-gradient that
(1.3) $\lim_{n\to\infty}\nabla^{g}F(y_{n})=0.$
Let $C_{3}\geq 0$ with $|F(x_{n})|\leq C_{3}$ for all $n\in\mathbb{N}$. Then $F(y_{n})=F(\phi(x_{n},t_{n}))\leq F(x_{n})\leq C_{3}$ for each $n\in\mathbb{N}$. By assumption, $F$ is bounded from below, so it follows that $(F(y_{n}))_{n\in\mathbb{N}}$ is bounded. Together with (1.3), the PS condition for $(F,g)$ implies that $(y_{n})_{n\in\mathbb{N}}$ has a subsequence $(y_{n_{k}})_{k\in\mathbb{N}}$ which converges to some $z\in\operatorname{Crit}F=\mathrm{Fix\;}\phi_{1}$. We want to show that $(x_{n_{k}})_{k\in\mathbb{N}}$ converges to $z$ as well. Let $C_{1},C_{2}\in(0,+\infty)$ be given as in Definition 1.5. For each $k\in\mathbb{N}$, $\gamma_{k}\colon[0,t_{n_{k}}]\to M$, $\gamma_{k}(t)=\phi(x_{n_{k}},t)$ is a $C^{1}$-path from $x_{n_{k}}$ to $y_{n_{k}}$, so it holds that
[…]
where we used property (ii) of a pseudo-gradient . We derive from property (iii) that
$\forall x\in M:\quad\|\nabla^{g}F(x)\|\leq C_{2}^{-1}\cdot DF_{x}(X(x))\quad\vee\quad\|\nabla^{g}F(x)\|^{2}\leq C_{2}^{-1}\cdot DF_{x}(X(x)),$
and thus $\|\nabla^{g}F(x)\|\leq C_{3}\left(DF_{x}(X(x))+(DF_{x}(X(x)))^{\frac{1}{2}}\right)$, where $C_{3}:=\max\{C_{2}^{-1},\sqrt{C_{2}^{-1}}\}$. Inserting this into the above computation shows that
[…]
where we have used Jensen’s inequality for concave functions. We derive from (1.2) that
$d_{M}(x_{n_{k}},y_{n_{k}})\leq C_{1}C_{3}\Big{(}\frac{1}{n_{k}}+\frac{1}{\sqrt{n_{k}}}\Big{)},$
which yields $\lim_{k\to\infty}d_{M}(x_{n_{k}},y_{n_{k}})=0$. Using the triangle inequality we derive that $(x_{n_{k}})_{k\in\mathbb{N}}$ converges to $z$ as well, which shows that $z\in\overline{A}\cap\mathrm{Fix\;}\phi_{1}$. Thus, $(\phi_{1},F)$ satisfies ($D_{2}$). ∎
2. Sequential subspace complexities as index functions
We want to define an index function via open covers of subsets by domains of continuous sequential motion planners. The latter are defined below and our index functions generalize the concept of relative complexity as introduced in [Far08, Section 4.3].
Throughout this section we choose and fix some $r\in\mathbb{N}$ with $r\geq 2$.
Given a path-connected topological space $X$, we consider the path fibration
$p_{r}\colon PX\to X^{r},\qquad p_{r}(\gamma)=\big{(}\gamma(0),\gamma\big{(}\tfrac{1}{r-1}\big{)},\dots,\gamma\big{(}\tfrac{r-2}{r-1}\big{)},\gamma(1)\big{)}.$
Definition 2.1.
Let $X$ be a path-connected topological space and let $A\subset X^{r}$.
  1. a)
    A sequential motion planner on $A$ is a local section $s\colon A\to PX$ of $p_{r}$.
  2. b)
    The $r$-th sequential subspace complexity of $A$, denoted by $\operatorname{\mathsf{TC}}_{r,X}(A)$, is given as the minimal value of $n\in\mathbb{N}$ for which there exist open subsets $U_{1},U_{2},\dots,U_{n}\subset X^{r}$ with $A\subset\bigcup_{i=1}^{n}U_{i}$, such that for each $i\in\{1,2,\dots,n\}$ there exists a continuous sequential motion planner $s_{i}\colon U_{i}\to PX$. In case there is no such $n$, we put $\operatorname{\mathsf{TC}}_{r,X}(A):=\infty$.
Our aim in this section is to show that under mild assumptions on $X$, we can define an index function on $X^{r}$ by putting
$\nu(A)=\operatorname{\mathsf{TC}}_{r,x}(A)$
for each $A\subset X^{r}$. For this purpose, we need to consider some preparatory technical statements.
The following lemma is an immediate consequence of the assertions of [Rud10, Remark 3.2] and an analogue of [Far06, Lemma 18.1]. For the reader’s convenience, we provide an explicit proof.
Lemma 2.2.
Let $X$ be a path-connected Hausdorff space, consider
$d\colon X^{r}\to X^{r},\quad d(x_{1},x_{2},\dots,x_{r}):=(x_{1},x_{1},\dots,x_{1}),$
and let $A\subset X^{r}$. The following statements are equivalent:
  1. (1)
    $A$ admits a continuous sequential motion planner.
  2. (2)
    There exists a deformation $h\colon A\times[0,1]\to X^{r}$ of $A$ with $h(a,1)=d(a)$ for each $a\in A$.
Proof.
(1) $\Rightarrow$ (2):  Let $s\colon A\to PX$ be a continuous sequential motion planner. Then
$h\colon A\times[0,1]\to X^{r},$
$h(a,t)=\Big{(}s(a)(0),s(a)\big{(}\tfrac{1-t}{r-1}\big{)},s(a)\big{(}\tfrac{2(1-t)}{r-1}\big{)},\dots,s(a)\big{(}\tfrac{(r-2)(1-t)}{r-1}\big{)},s(a)(1-t)\Big{)},$
is a deformation with the desired properties.
(2) $\Rightarrow$ (1):  Let $h\colon A\times[0,1]\to X^{r}$ be a deformation of $A$ with $h(a,1)=d(a)$ for all $a\in A$. Denote the components of $h$ by $h(a,t)=(h_{0}(a,t),\dots,h_{r-1}(a,t))$ for $a\in A$ and $t\in[0,1]$. Define a map $s\colon A\to PX$ by
(2.1) $(s(a))(t)=\begin{cases}h_{j-1}(a,2((r-1)t-j+1))&\text{if }t\in\Big{[}\frac{j-1}{r-1},\frac{j-\frac{1}{2}}{r-1}\Big{)},\\ h_{j}(a,2(j-(r-1)t))&\text{if }t\in\Big{[}\frac{j-\frac{1}{2}}{r-1},\frac{j}{r-1}\Big{]},\end{cases}\qquad\forall j\in\{1,2,\dots,r-1\}.$
One checks without difficulties that $s(a)$ is indeed a continuous path with $(p_{r}\circ s)(a)=a$ for each $a\in A$. Using that $h$ is continuous and $X$ is Hausdorff, one derives from [Bre93, Theorem VII.2.4] that $s$ is continuous. ∎
To establish the continuity property of $\nu(A)=\operatorname{\mathsf{TC}}_{r,X}(A)$, we need to make an additional assumption on the space under consideration. To clarify the type of spaces that we will consider, we recall some terminology.
Definition 2.3.
A topological space $Y$ is called an absolute neighborhood retract (ANR) if it is metrizable and satisfies the following condition:
Let $X$ be a metrizable space and let $A\subset X$ be closed. Then for any continuous map $f\colon A\to Y$ there are an open neighborhood $U\subset X$ of $A$ and a continuous map $F\colon U\to Y$ with $F|_{A}=f$.
Remark 2.4.
Note that this definition coincides with the one used in [CLOT03] and [Pal66a], but differs from the one in [Hu65], which we will occasionally use as a reference. More precisely, our notion of an ANR coincides with the definition of an absolute neighborhood extensor (ANE) in [Hu65].
The following lemma is an on-the-nose analogue of [Mes21, Lemma 1.6] which considers spherical complexities and whose proof is a straightforward analogue of the corresponding statement for Lusternik-Schnirelmann category, see [CLOT03, Lemma 1.11].
Lemma 2.5.
Let $X$ be a path-connected ANR and let $A\subset X^{r}$ be a closed subset which admits a continuous sequential motion planner. Then there exists an open neighborhood $U\subset X^{r}$ of $A$ which admits a continuous sequential motion planner.
Proof.
Let again $d\colon X^{r}\to X^{r}$, $d(x_{1},x_{2},\dots,x_{r})=(x_{1},x_{1},\dots,x_{1})$. By Lemma 2.2.a), there exists a deformation $h\colon A\times[0,1]\to X^{r}$ with $h(a,1)=d(a)$ for each $a\in A$. Put $Q:=(X^{r}\times\{0,1\})\cup(A\times[0,1])$ and define
(2.2) $H\colon Q\to X^{r},\qquad H(x,t)=\begin{cases}x&\text{if }t=0,\\ d(x)&\text{if }t=1,\\ h(x,t)&\text{if }x\in A.\end{cases}$
One checks without difficulties that $H$ is well-defined and continuous. Since $X$ is an ANR, $X^{r}$ is an ANR as well. Hence, there exist an open neighborhood $V\subset X^{r}\times[0,1]$ of $Q$ and a continuous map $K\colon V\to X^{r}$ with $K|_{Q}=H$. Since $[0,1]$ is compact, there exists an open subset $U\subset X^{r}$ with $U\times[0,1]\subset V$ and $A\subset U$. By definition of the map, $K|_{U\times[0,1]}$ is a deformation of $U$ with $K(x,1)=d(x)$ for each $x\in U$. Since $X$ is by assumption metrizable, it is a Hausdorff space, so it follows from Lemma 2.2.a) that $U$ admits a continuous sequential motion planner. ∎
Using the previous lemma, we can now establish that sequential subspace complexities indeed form index functions.
Proposition 2.6.
Let $X$ be path-connected.
  1. a)
    For all $A\subset B\subset X^{r}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(A)\leq\operatorname{\mathsf{TC}}_{r,X}(B).$
  2. b)
    For all $A,B\subset X^{r}$ it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,X}(A)+\operatorname{\mathsf{TC}}_{r,X}(B).$
  3. c)
    Let $A\subset X^{r}$ be a closed subset. Then there exists an open subset $U\subset X^{r}$ with $A\subset U$, such that
    $\operatorname{\mathsf{TC}}_{r,X}(A)=\operatorname{\mathsf{TC}}_{r,X}(U).$
  4. d)
    Assume that $X$ is normal. Let $A,B\subset X^{r}$ be closed with $A\cap B=\emptyset$. Then
    $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)=\max\{\operatorname{\mathsf{TC}}_{r,X}(A),\operatorname{\mathsf{TC}}_{r,X}(B)\}.$
Proof.
  1. a)
    This is obvious.
  2. b)
    This follows immediately from the definition of $\operatorname{\mathsf{TC}}_{r,X}$.
  3. c)
    Let $k:=\operatorname{\mathsf{TC}}_{r,X}(A)$. Then by definition of $\operatorname{\mathsf{TC}}_{r,X}(A)$ there are $k$ open domains $U_{1},\dots,U_{k}\subset X^{r}$ of continuous sequential motion planners, such that $A\subset\bigcup_{i=1}^{k}U_{i}$. Put $U:=\bigcup_{i=1}^{k}U_{i}$. Since $A\subset U$, we derive from a) that $k\leq\operatorname{\mathsf{TC}}_{r,X}(U)$. However, since $U_{1},\dots,U_{k}$ form an open cover of $U$ for which each element admits a continuous sequential motion planner, it follows that $\operatorname{\mathsf{TC}}_{r,X}(U)\leq k$. Together, the two inequalities show the claim.
  4. d)
    It follows immediately from a) that $\max\{\operatorname{\mathsf{TC}}_{r,X}(A),\operatorname{\mathsf{TC}}_{r,X}(B)\}\leq\operatorname{\mathsf{TC}}_{r,X}(A\cup B)$, so it only remains to show the opposite inequality. Let $m:=\operatorname{\mathsf{TC}}_{r,X}(A)$ and $n:=\operatorname{\mathsf{TC}}_{r,X}(B)$ and assume w.l.o.g. that $m\geq n$. Let $U_{1},\dots,U_{m},V_{1},\dots,V_{n}\subset X^{r}$ be open subsets, such that there are continuous sequential motion planners $s_{i}\colon U_{i}\to PX$ and $t_{j}\colon V_{j}\to PX$ for all $i\in\{1,2,\dots,m\}$ and $j\in\{1,2,\dots,n\}$ and such that
    $A\subset\bigcup_{i=1}^{m}U_{i},\qquad B\subset\bigcup_{i=1}^{n}V_{j}.$
    If $m>n$, then put $V_{j}:=\emptyset$ for all $j\in\{m+1,\dots,n\}$. Since $X$ is normal, $X^{r}$ is normal as well, so there are open subsets $U_{0},V_{0}\subset X^{r}$ with $A\subset U_{0}$, $B\subset V_{0}$ and $U_{0}\cap V_{0}=\emptyset$. Put $U_{i}^{\prime}:=U_{0}\cap U_{i}$ and $V_{i}^{\prime}:=V_{0}\cap V_{i}$ for all $i\in\{1,2,\dots,m\}$. Then for each $i$ we obtain a continuous sequential motion planner
    $U^{\prime}_{i}\cup V_{i}^{\prime}\to PX,\qquad x\mapsto\begin{cases}s_{i}(x)&\text{if }x\in U_{i}^{\prime},\\ t_{i}(x)&\text{if }x\in V_{i}^{\prime}.\end{cases}$
    Since $A\cup B\subset\bigcup_{i=1}^{m}(U_{i}^{\prime}\cup V_{i}^{\prime})$, this shows that $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)\leq m=\max\{\operatorname{\mathsf{TC}}_{r,X}(A),\operatorname{\mathsf{TC}}_{r,X}(B)\}$.
We next want to show that under mild assumptions on $X$ the index function $\operatorname{\mathsf{TC}}_{r,X}$ is indeed supervariant with respect to time-1 maps of deformations.
Proposition 2.7.
Let $X$ be a path-connected ANR, let $A\subset X^{r}$ be closed and let $\Phi\colon A\times[0,1]\to X^{r}$ be a deformation of $A$. Put $\Phi_{1}:=\Phi(\cdot,1)\colon A\to X^{r}$. If $\Phi_{1}(A)$ is closed, then
$\operatorname{\mathsf{TC}}_{r,X}(\Phi_{1}(A))\geq\operatorname{\mathsf{TC}}_{r,X}(A).$
Proof.
Put $k:=\operatorname{\mathsf{TC}}_{r,X}(\Phi_{1}(A))$ and let $U_{1},\dots,U_{k}\subset X^{r}$ be open subsets with $\Phi_{1}(A)\subset\bigcup_{j=1}^{k}U_{j}$, such that each $U_{j}$ admits a continuous sequential motion planner $s_{j}\colon U_{j}\to PX$. Since $X$ is an ANR, $X$ is metrizable, hence normal. As a closed subspace of $X^{r}$, $\Phi_{1}(A)$ is normal as well. Thus, see e.g. [CLOT03, Theorem A.1.(6)], we can find closed subsets $B_{j}\subset U_{j}$ for $j\in\{1,2,\dots,k\}$ with $\Phi_{1}(A)=\bigcup_{j=1}^{k}B_{j}$. Put $B_{j}^{\prime}:=\Phi_{1}^{-1}(B_{j})$ for each $j\in\{1,2,\dots,k\}$. Then $A\subset\Phi_{1}^{-1}(\Phi_{1}(A))\subset\bigcup_{j=1}^{k}B_{j}^{\prime}$. Moreover, for each $j$ we obtain a continuous motion planner $s_{j}^{\prime}\colon B_{j}^{\prime}\to PX$ as follows: Denote the components of $\Phi$ by $\phi_{j}\colon A\times[0,1]\to X$, $j\in\{1,2,\dots,r\}$, such that $\Phi(x,t)=(\phi_{1}(x,t),\dots,\phi_{r}(x,t))$. Given $x=(x_{0},x_{1},\dots,x_{r-1})\in B_{j}^{\prime}$, $j\in\{1,2,\dots,r-1\}$ and $t\in\left[\frac{j-1}{r-1},\frac{j}{r-1}\right]$, we put
(2.3) $(s_{j}^{\prime}(x))(t):=\begin{cases}\phi_{j-1}(x,3(r-1)t-3j+3)&\text{if }t\in\Big{[}\frac{j-1}{r-1},\frac{j-\frac{2}{3}}{r-1}\Big{)},\\ (s_{j}(\Phi_{1}(x)))\left(3t+\frac{1-2j}{r-1}\right)&\text{if }t\in\Big{[}\frac{j-\frac{2}{3}}{r-1},\frac{j-\frac{1}{3}}{r-1}\Big{)},\\ \phi_{j}(x,3j-3(r-1)t)&\text{if }t\in\Big{[}\frac{j-\frac{1}{3}}{r-1},\frac{j}{r-1}\Big{]}.\end{cases}$
One checks that $s_{j}^{\prime}$ is continuous and that $(p_{r}\circ s_{j}^{\prime})(x)=x$ for each $x\in B_{j}^{\prime}$, so each $s_{j}^{\prime}$ is a continuous sequential motion planner. Since $B_{j}^{\prime}=\Phi_{1}^{-1}(B_{j})$ is closed and since $X^{r}$ is an ANR, it follows from Lemma 2.5 that there is an open neighborhood $V_{j}$ of $B_{j}^{\prime}$ which admits a continuous sequential motion planner. Since $A\subset\bigcup_{j=1}^{k}V_{j}$, this shows that $\operatorname{\mathsf{TC}}_{r,X}(A)\leq k$. ∎
Note that the condition on $\Phi_{1}(A)$ to be closed in Proposition 2.7 will be satisfied for every closed subset if $\Phi_{1}$ is a homeomorphism or if $X$ is compact.
Corollary 2.8.
Let $X$ be a path-connected ANR and let $\varphi\colon X^{r}\to X^{r}$ be a closed map which is homotopic to the identity. Then $\operatorname{\mathsf{TC}}_{r,X}\colon\mathfrak{P}(X^{r})\to\mathbb{N}_{0}\cup\{\infty\}$ is a $\varphi$-supervariant index function.
Proof.
It follows immediately from Proposition 2.6 that $\operatorname{\mathsf{TC}}_{r,X}\colon\mathfrak{P}(X^{r})\to\mathbb{N}_{0}\cup\{\infty\}$ is an index function. Let $H\colon X^{r}\times[0,1]\to X^{r}$ be a homotopy from $\operatorname{id}_{X^{r}}$ to $\varphi$. Then $H|_{A\times[0,1]}$ is a deformation of $A$ for every $A\subset X^{r}$. In particular, if $A$ is closed, it follows from Proposition 2.7 that
$\operatorname{\mathsf{TC}}_{r,X}(\varphi(A))=\operatorname{\mathsf{TC}}_{r,X}(H(A\times\{1\}))\geq\operatorname{\mathsf{TC}}_{r,X}(A).$
Thus, $\operatorname{\mathsf{TC}}_{r,X}$ is $\varphi$-supervariant. ∎
After establishing the index function properties, we next want to apply Theorem 1.4 to sequential subspace complexities in several settings. Before doing so, let us introduce some shorthand notation that will be useful in the statements of several of the following results.
Definition 2.9.
  1. a)
    Let $X$ be a topological space, $F\colon X\to\mathbb{R}$ and $\varphi\colon X\to X$. For each $\mu\in\mathbb{R}$ we put
    $\mathcal{C}_{\mu}(\varphi,F):=\pi_{0}(F_{\mu}\cap\mathrm{Fix\;}\varphi),$
    i.e. the set of connected components of $F_{\mu}\cap\mathrm{Fix\;}\varphi$.
  2. b)
    Let $M$ be a smooth Banach manifold and $F\in C^{1}(M)$. For each $\mu\in\mathbb{R}$ we put
    $\mathcal{C}_{\mu}(F):=\pi_{0}(F_{\mu}\cap\operatorname{Crit}F).$
By Corollary 2.8 we can apply Theorem 1.4 to the index function $\operatorname{\mathsf{TC}}_{r,X}(\cdot)$, to obtain the following result.
Theorem 2.10.
Let $X$ be a path-connected ANR, let $\varphi\colon X^{r}\to X^{r}$ be continuous and homotopic to the identity and let $F\colon X^{r}\to\mathbb{R}$ be a Lyapunov function for $\varphi$ which is bounded from below. Assume that $F(\mathrm{Fix\;}\varphi)$ is discrete and that $(\varphi,F)$ satisfies condition (D).
  1. a)
    For all $\lambda\in\mathbb{R}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,X}(F_{\mu}\cap\mathrm{Fix\;}\varphi).$
  2. b)
    If $\lambda\in\mathbb{R}$ is such that $\mathcal{C}_{\mu}(\varphi,F)$ is finite for all $\mu\leq\lambda$, then
    $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(\varphi,F)\}.$
Proof.
  1. a)
    This is an immediate consequence of combining Corollary 2.8 and Theorem 1.4.
  2. b)
    Let $\mu\in(-\infty,\lambda]$. Since $X$ is a Hausdorff space, $\mathrm{Fix\;}\varphi$ is a closed subset of $X^{r}$. Hence, for every $\mu\in\mathbb{R}$, the subset $F_{\mu}\cap\mathrm{Fix\;}\varphi$ is closed in $X^{r}$ and since every $C\in\mathcal{C}_{\mu}(\varphi,F)$ is closed in $F_{\mu}\cap\mathrm{Fix\;}\varphi$, it is closed in $X^{r}$ as well. Since each $\mathcal{C}_{\mu}(\varphi,F)$ is finite, we can iteratively apply Proposition 2.6.d) and obtain
    $\operatorname{\mathsf{TC}}_{r,X}(F_{\mu}\cap\mathrm{Fix\;}\varphi)=\operatorname{\mathsf{TC}}_{r,X}\Big{(}\bigcup_{C\in\mathcal{C}_{\mu}(\varphi,F)}C\Big{)}=\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}.$
    Combining this observation with part a) shows the claim.
Since every paracompact Banach manifold is an ANR by [Pal66a, Theorem 5] and since every Riemannian Banach manifold is paracompact, we obtain the following as a direct consequence of Theorem 2.10.
Corollary 2.11.
Let $(X,g)$ be a connected and complete Riemannian Banach manifold and let $F\in C^{1}(X^{r})$ be bounded from below. Assume that $(F,g^{r})$ satisfies the PS condition, where $g^{r}$ denotes the product metric of $g$ on $X^{r}$, and that $F(\operatorname{Crit}F)$ is discrete. Then
$\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}\qquad\forall\lambda\in\mathbb{R}.$
Proof.
By [Pal66b, Theorem 4.4], see also [Str08, Lemma II.3.9], the function $F$ admits a pseudo-gradient $V\colon X^{r}\to TX^{r}$ in the sense of Definition 1.5. Let $\phi_{1}\colon X^{r}\to X^{r}$ be the time-1 map of the flow of $-V$, which is well-defined by the completeness of $(X^{r},g^{r})$. The flow yields a homotopy from $\phi_{1}$ to the identity map. By Theorem 1.10, $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D). Moreover, by Lemma 1.9, $\mathcal{C}_{\mu}(F)$ is finite for all $\mu\in\mathbb{R}$. Thus, the claim follows directly from Theorem 2.10.b), since by definition of $\phi_{1}$ and by Remark 1.6.(3) it holds that $\mathrm{Fix\;}\phi_{1}=\mathrm{sing}\,V=\operatorname{Crit}F$. ∎
3. Navigation functions and sequential subspace complexities
In this section, we want to study consequences of Corollary 2.11 a for special types of differentiable functions, namely $r$-navigation functions. These are straightforward generalizations of the notion of a navigation function as introduced in [Far08, Chapter 4], see also [Mes24, Section 4.1] for an exposition.
Definition 3.1.
Let $r\in\mathbb{N}$ with $r\geq 2$ and let $X$ be a connected paracompact Banach manifold. We call a continuously differentiable function $F\colon X^{r}\to\mathbb{R}$ an $r$-navigation function if
  1. (i)
    $F$ satisfies the PS condition with respect to a complete Riemannian metric on $X^{r}$,
  2. (ii)
    $F(x)\geq 0$ for all $x\in X^{r}$ and $F^{-1}(\{0\})=\Delta_{r,X}$.
Remark 3.2.
Let $r\in\mathbb{N}$ with $r\geq 2$
  1. (1)
    By definition of an $r$-navigation function $F\colon X^{r}\to\mathbb{R}$, its global minimum is 00 and it holds that
    $\operatorname{Crit}_{0}F=F_{0}=\Delta_{r,X}.$
  2. (2)
    If $X$ is a finite-dimensional closed manifold, then any $F\in C^{1}(X^{r})$ which satisfies property (ii) of Definition 3.1 is an $r$-navigation function, since any $C^{1}$-function on a closed manifold satisfies the PS condition with respect to any Riemannian metric.
It follows from Remark 3.2.(1) that if $X$ is not a point, then any $r$-navigation function $F\colon X^{r}\to\mathbb{R}$ will have infinitely many critical points, since $\Delta_{r,X}\subset\operatorname{Crit}F$. However, we can use subspace complexities to estimate the numbers of critical points whose value is positive as the following application of Corollary 2.11 shows.
Theorem 3.3.
Let $r\in\mathbb{N}$ with $r\geq 2$, let $X$ be a connected paracompact Banach manifold and let $F\colon X^{r}\to\mathbb{R}$ be an $r$-navigation function. Then
$\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1\qquad\forall\lambda\in\mathbb{R}.$
Proof.
Since $F$ is non-negative, we obtain from Corollary 2.11 and Remark 3.2.(1) that
$\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})$ $\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+\operatorname{\mathsf{TC}}_{r,X}(\Delta_{r,X})$
$=\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,X}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1$
for each $\lambda\in(0,+\infty)$. Here, we used that $s\colon\Delta_{r,X}\to PX$, $s(x,\dots,x)=c_{x}$, where $c_{x}$ denotes the constant path in $x$, is a continuous sequential motion planner which extends to an open neighborhood of $\Delta_{r,X}$ by Lemma 2.5, so that $\operatorname{\mathsf{TC}}_{r,X}(\Delta_{r,X})=1$. ∎
Note that Theorem 3.3 is a straightforward generalization of [Far08, Theorem 4.32], which deals with the case of $r=2$ and a smaller class of functions. The following example is discussed in [Far08, Example 4.31] in the case of $r=2$ and under the additional assumption that it is a Morse-Bott function.
Example 3.4.
[…]
$F_{r}\colon M^{r}\to\mathbb{R},\qquad F_{r}(x_{1},\dots,x_{r})=\sum_{i=1}^{r-1}\|x_{i}-x_{i+1}\|^{2},$
is a smooth function, whose differential is straightforwardly computed as
(3.1) $(DF_{r})_{(x_{1},\dots,x_{r})}(v_{1},\dots,v_{r})=2\sum_{i=1}^{r-1}\left<x_{i}-x_{i+1},v_{i}-v_{i+1}\right>$
with respect to the Euclidean inner product. Evidently, $F$ is non-negative and one checks that
$F_{r}(x_{1},x_{2},\dots,x_{r})=0\quad\Leftrightarrow\quad x_{1}=x_{2}=\dots=x_{r},$
i.e. that $F_{r}^{-1}(\{0\})=\Delta_{r,M}$. Thus, $F_{r}$ has property (ii) of Definition 3.1. Since $M$ is compact, this yields by Remark 3.2.(2) that $F_{r}$ is an $r$-navigation function.
Studying $2$-navigation functions of this particular type leads to an interesting geometric consequence for hypersurfaces in Euclidean space.
Theorem 3.5.
Let $n\in\mathbb{N}$ and let $M\subset\mathbb{R}^{n}$ be a compact hypersurface and let
$\alpha(M):=\#\left\{\{x,y\}\subset M\ \middle|\ x\neq y,\ T_{x}M=T_{y}M=(x-y)^{\perp}\right\}.$
Then
$\alpha(M)\geq\operatorname{\mathsf{TC}}(M)-1.$
In particular, there are at least $\operatorname{\mathsf{TC}}(M)-1$ distinct unordered pairs of points $\{x,y\}\subset M$ which satisfy $T_{x}M=T_{y}M$.
Proof.
[…]
(3.2) $\operatorname{\mathsf{TC}}(M)=\operatorname{\mathsf{TC}}(F^{\lambda})\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1.$
A straightforward computation further shows that
$DF_{(x,y)}(v,w)=2\left<x-y,v-w\right>\quad\forall x,y\in M,\,v\in T_{x}M,\,w\in T_{y}M,$
from which it follows that
$(x,y)\in\operatorname{Crit}F\quad\Leftrightarrow\quad v\in(x-y)^{\perp}\quad\forall v\in T_{x}M\cup T_{y}M.$
So for each $(x,y)\in\operatorname{Crit}F$ it holds that $T_{x}M\subset(x-y)^{\perp}$ and $T_{y}M\subset(x-y)^{\perp}$. Since $M$ is a hypersurface, it holds that $\dim T_{x}M=\dim T_{y}M=\dim(x-y)^{\perp}=n-1$, so that these three spaces have to coincide. In particular, since $F$ is a $2$-navigation function, this shows that
$\alpha(M)=\#\{(x,y)\in\operatorname{Crit}F\mid x\neq y\}=\#\{(x,y)\in\operatorname{Crit}F\mid F(x,y)>0\}.$
Since $\alpha(M)$ is by assumption finite, $F_{\mu}\cap\operatorname{Crit}F$ is finite for each $\mu>0$. Thus, each $C\in\mathcal{C}_{\mu}(F)$ consists of a single point, which immediately yields $\operatorname{\mathsf{TC}}_{M}(C)=1$. We derive from (3.2) that
$\operatorname{\mathsf{TC}}(M)$$\leq\#\{\mu\in(0,\lambda]\mid F_{\mu}\cap\operatorname{Crit}F\neq\emptyset\}+1$
#$\leq\#\{(x,y)\in\operatorname{Crit}F\mid F(x,y)>0\}+1=\alpha(M)+1.$
The claim immediately follows. ∎
So far, our guideline was to estimate the number of critial points of functions from below in terms of sequential subspace complexities. In the following, we want to go the opposite way and show in an example how to use $r$-navigation functions to estimates values of sequential TCs from above.
Let $k,r\in\mathbb{N}$ with $r\geq 2$ and choose numbers $n_{1},n_{2},\ldots,n_{k}\in\mathbb{N}$. The sequential topological complexities of
(3.3) $M=\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\ldots\times\mathbb{S}^{n_{k}},$
were originally computed by Basabe, González, Rudyak and Tamaki in [BGRT14, Corollary 3.12]. In particular, they have shown that if all of $n_{1},n_{2},\dots,n_{k}$ are odd, then
$\operatorname{\mathsf{TC}}_{r}(M)=k(r-1)+1.$
They obtained this value as a lower bound for $\operatorname{\mathsf{TC}}_{r}(M)$ by a standard cohomological computation and as an upper bound using an inequality for sequential TCs of product spaces in terms of the sequential TCs of its factors. Here, we want to show that the upper bound can also be obtained by using $r$-navigation functions.
For $n\in\mathbb{N}$ we consider the $n$-sphere as
$\mathbb{S}^{n}=\{x\in\mathbb{R}^{n+1}\mid\|x\|=1\},$
[…]
$f_{n}\colon(\mathbb{S}^{n})^{r}\to\mathbb{R},\qquad f_{n}(x_{1},\dots,x_{r})=\sum_{i=1}^{r-1}\|x_{i}-x_{i+1}\|^{2},$
from Example 3.4.
Lemma 3.6.
For each $n\in\mathbb{N}$ it holds that
$\operatorname{Crit}f_{n}=\big{\{}(x_{1},\dots,x_{r})\in(\mathbb{S}^{n})^{r}\ \big{|}\ x_{i}\in\{x_{1},-x_{1}\}\ \forall i\in\{2,3,\dots,r\}\big{\}}.$
Proof.
Let $x=(x_{1},\dots,x_{r})\in\operatorname{Crit}f_{n}$. We will prove by induction over $i$ that $\{x_{1},x_{i}\}$ is linearly dependent for each $i\in\{2,3,\dots,r\}$. Using (3.1), we first compute for each $v_{1}\in T_{x_{1}}\mathbb{S}^{n}$ that
$0=DF_{x}(v_{1},0,\dots,0)=2\left<x_{1}-x_{2},v_{1}\right>=2\left<x_{2},v_{1}\right>.$
Thus, $T_{x_{1}}\mathbb{S}^{n}\subset\{x_{2}\}^{\perp}$ and since $T_{x_{1}}\mathbb{S}^{n}=\{x_{1}\}^{\perp}$ it follows for dimensional reasons that $\{x_{1}\}^{\perp}=\{x_{2}\}^{\perp}$, so $x_{1}$ and $x_{2}$ are linearly dependent.
Assume as induction hypothesis that $x_{1}$ and $x_{j}$ are linearly dependent for all $j\in\{2,3,\dots,i\}$ for some $i\geq 2$. We compute for each $v_{i}\in T_{x_{i}}\mathbb{S}^{n}$ that
[…]
since $x_{i},v_{i}\right>=0$. In analogy with the case of $i=2$, this shows that $T_{x_{i}}\mathbb{S}^{n}\subset\{x_{i-1}+x_{i+1}\}^{\perp}$ and consequently that $x_{i}$ and $x_{i-1}+x_{i+1}$ are linearly dependent, i.e. there are $\lambda_{1},\lambda_{2}\in\mathbb{R}$ with $(\lambda_{1},\lambda_{2})\neq(0,0)$ and
$\lambda_{1}x_{i}+\lambda_{2}(x_{i-1}+x_{i+1})=0.$
If $\lambda_{2}=0$, then it would immediately follow that $x_{i}=0$, which is a contradiction. Hence, $\lambda_{2}\neq 0$ and we can rewrite the equation as $x_{i+1}=-\frac{\lambda_{1}}{\lambda_{2}}x_{i}-x_{i-1}$. By our induction hypothesis, both $x_{i}$ and $x_{i-1}$ are linearly dependent of $x_{1}$, so it follows immediately that $x_{i+1}$ is linearly dependent of $x_{1}$ as well, which completes our induction. Hence, we have shown that $x_{i}\in\mathbb{R}x_{1}$ for each $i$ and since $x_{1},\dots,x_{r}\in\mathbb{S}^{n}$, this shows that $x_{i}\in\{x_{1},-x_{1}\}$ for each $i\in\{1,2,\dots,r\}$. Conversely, a straightforward calculation which we omit shows that any such point is indeed a critical point of $f_{n}$. ∎
Let $n\in\mathbb{N}$ be arbitrary. Using Lemma 3.6 one computes for each $(x_{1},\dots,x_{r})\in\operatorname{Crit}f_{n}$ that
$f_{n}(x_{1},\dots,x_{r})=4m,\quad\text{where }\ \ m=\#\{i\in\{2,3,\dots,r\}\mid x_{i}\neq x_{1}\}.$
Thus, the set of critical values of $f_{n}$ is given as
$f_{n}(\operatorname{Crit}f_{n})=\{0,4,8,\dots,4(r-1)\}.$
For $n_{1},n_{2},\dots,n_{k}\in\mathbb{N}$ we let again $M=\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\mathbb{S}^{n_{k}}$ and consider the function
$F\colon M^{r}\to\mathbb{R},\quad F((x_{11},\ldots,x_{1k}),\ldots,(x_{r1},\ldots,x_{rk}))=\sum_{j=1}^{k}f_{n_{j}}(x_{1j},x_{2j},\dots,x_{rj}).$
Since the $f_{n_{j}}$ are $r$-navigation functions, it follows that $F$ is non-negative with $F^{-1}(\{0\})=\Delta_{r,M}$, so by Remark 3.2.(2), $F$ is an $r$-navigation function as well. Moreover, it follows that
$\operatorname{Crit}F$ =$=\left\{\left((x_{11},\dots,x_{1k}),\dots,(x_{r1},\dots,x_{rk})\right)\in M^{r}\ \middle|\ (x_{1j},\dots,x_{rj})\in\operatorname{Crit}f_{n_{j}}\ \forall j\in\{1,2,\dots,k\}\right\}$
=$=\left\{\left((x_{11},\dots,x_{1k}),\dots,(x_{r1},\dots,x_{rk})\right)\in M^{r}\ \middle|\ x_{ij}\in\{x_{1j},-x_{1j}\}\ \forall i\in\{1,\dots,r\},\ j\in\{1,\dots,k\}\right\},$
where we have used Lemma 3.6. The critical values of $F$ are further given by
$F(\operatorname{Crit}F)=\Big{\{}\sum_{j=1}^{k}\mu_{j}\ \Big{|}\ \mu_{j}\in f_{n_{j}}(\operatorname{Crit}f_{n_{j}})\ \forall j\in\{1,2,\dots,k\}\Big{\}}=\{0,4,8,\dots,4k(r-1)\}.$
Therefore, it follows from Theorem 3.3 that
(3.4) $\operatorname{\mathsf{TC}}_{r}(M)\leq\sum_{i=1}^{k(r-1)}\max\{\operatorname{\mathsf{TC}}_{r,M}(\Sigma)\mid\Sigma\in\mathcal{C}_{4i}(F)\}+1.$
Assume for the rest of this section that $n_{1},\dots,n_{k}\in\mathbb{N}$ are odd numbers.
Let $i\in\{0,1,\dots,k(r-1)\}$ and let $\Sigma\in\mathcal{C}_{4i}(F)$. By the above description of $\operatorname{Crit}F$, there are numbers $\lambda_{j,\ell}\in\{-1,1\}$, where $j\in\{1,2,\dots,k\}$ and $\ell\in\{2,3,\dots,r\}$, such that
$\Sigma=\{((x_{1},\ldots,x_{k}),(\lambda_{1,2}x_{1},\ldots,\lambda_{k,2}x_{k}),\ldots,(\lambda_{1,r}x_{1},\ldots,\lambda_{k,r}x_{k}))\mid(x_{1},\ldots,x_{k})\in M\}.$
For each $j\in\{1,2,\ldots,k\}$ we further put $\lambda_{j,1}:=1$, choose a nowhere-vanishing continuous vector field $Y_{j}$ on $\mathbb{S}^{n_{j}}$ and normalize it by putting
$X_{j}(x):=\frac{(r-1)\pi}{\|Y_{j}(x)\|}\cdot Y_{j}(x)\qquad\forall x\in\mathbb{S}^{n_{j}},\ j\in\{1,2,\dots,k\}.$
Note that by this normalization, $X_{j}(x_{j})$ is the tangent vector of a geodesic segment $[a,a+\frac{1}{r-1}]\to\mathbb{S}^{n_{j}}$ from $x_{j}$ to $-x_{j}$ for all $a\in\mathbb{R}$, $x_{j}\in\mathbb{S}^{n_{j}}$, $j\in\{1,2,\dots,k\}$. We define a sequential motion planner $s\colon\Sigma\to PM$ as follows:
$s((x_{1},\ldots,x_{k}),(\lambda_{1,2}x_{1},\ldots,\lambda_{k,2}x_{k}),\ldots,(\lambda_{1,r}x_{1},\ldots,\lambda_{k,r}x_{k}))=:(\gamma_{1},\dots,\gamma_{k}).$
Here, for $j\in\{1,2,\dots,k\}$ we let
$\gamma_{j}\colon[0,1]\to\mathbb{S}^{n_{j}},\qquad\gamma_{j}=\gamma_{j1}*\gamma_{j2}*\dots*\gamma_{j(r-1)},$
where * denotes the concatenation of paths, and where $\gamma_{j\ell}\colon[\frac{\ell-1}{r-1},\frac{\ell}{r-1}]\to\mathbb{S}^{n_{j}}$ is for each $\ell\in\{1,2,\dots,r-1\}$ given as
  • the constant path in $x_{j}$ if $\lambda_{j,\ell}=\lambda_{j,(\ell+1)}=1$,
  • the constant path in $-x_{j}$ if $\lambda_{j,\ell}=\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $x_{j}$ to $-x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(x_{j})$ if $\lambda_{j,\ell}=1$ and $\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $-x_{j}$ to $x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(-x_{j})$ if $\lambda_{j,\ell}=-1$ and $\lambda_{j,(\ell+1)}=1$.
One checks that the map $s$ we obtain in this way is indeed a continuous sequential motion planner on $\Sigma$. By Lemma 2.5 extends to a continuous sequential motion planner on an open neighborhood of $\Sigma$. Thus,
$\operatorname{\mathsf{TC}}_{r,M}(\Sigma)=1\quad\forall\Sigma\in\mathcal{C}_{4i}(F),\ i\in\{0,1,\dots,k(r-1)\}.$
Inserting this into (3.4) shows that
$\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\times\mathbb{S}^{n_{k}})\leq k(r-1)+1$
for all odd numbers $n_{1},n_{2},\dots,n_{k}\in\mathbb{N}$.
4. Sequential topological complexity of sublevel sets
In this section, we want to derive upper bounds on the sequential TC of a sublevel set of a function $f\colon X\to\mathbb{R}$ from the results of section 2. More precisely, we will apply those results to functions $F\colon X^{r}\to\mathbb{R}$ of the form
$F(x_{1},\dots,x_{r})=f(x_{1})+f(x_{2})+\dots+f(x_{r})$
and will rephrase the upper bounds in terms of $f$. While we have previously studied subspace complexities of the form $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})$, it turns out that we can find upper bounds for $\operatorname{\mathsf{TC}}_{r}(f^{\lambda})$, i.e. for the sequential TCs of $f^{\lambda}$ and not just for sequential subspace complexities.
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$.
We begin by making some technical preparations on condition (D) for product functions.
Lemma 4.1.
Let $X$ be a topological space, let $\varphi\colon X\to X$ be continuous and let $f\colon X\to\mathbb{R}$ be a Lyapunov function for $\varphi$ which is bounded from below. Consider
$\Phi\colon X^{r}\to X^{r},\quad\Phi(x_{1},x_{2},\dots,x_{r}):=(\varphi(x_{1}),\varphi(x_{2}),\dots,\varphi(x_{r})),$
$F\colon X^{r}\to\mathbb{R},\quad\ \ F(x_{1},x_{2},\dots,x_{r}):=f(x_{1})+f(x_{2})+\dots+f(x_{r}).$
Assume that $(\varphi,f)$ satisfies condition (D).
  1. a)
    Then $(\Phi,F)$ satisfies condition (D).
  2. b)
    Let $A\subset X^{r}$ satisfy $\Phi(A)\subset A$ and
    (4.1) $\overline{A}\cap\mathrm{Fix\;}\Phi\subset A.$
    Then $(\Phi|_{A},F|_{A})$ satisfies condition (D).
Proof.
  1. a)
    We begin by considering condition ($D_{1}$). Let $x=(x_{1},\dots,x_{r})\in X^{r}$ with $\Phi(x)\neq x$. Then there exists $i\in\{1,2,\dots,r\}$ with $\varphi(x_{i})\neq x_{i}$. Since $(\varphi,f)$ satisfies ($D_{1}$), this yields $f(\varphi(x_{i}))<f(x_{i})$. Since $f$ is a Lyapunov function for $\varphi$, we derive that
    $F(\Phi(x))=\sum_{j=1}^{r}f(\varphi(x_{j}))<\sum_{j=1}^{r}f(x_{j})=F(x).$
    Thus, $(\Phi,F)$ satisfies ($D_{1}$).
    Concerning ($D_{2}$), we let $B\subset X^{r}$ be such that $\inf\{F(x)-F(\Phi(x))\mid x\in B\}=0$ and such that $F$ is bounded on $B$. Let $(x_{n})_{n\in\mathbb{N}}$ be a sequence in $B$ with
    $\lim_{n\to\infty}(F(x_{n})-F(\Phi(x_{n})))=0.$
    We denote the components of the $x_{n}$ by $x_{n}=(x_{1,n},\dots,x_{r,n})$ for each $n\in\mathbb{N}$. Since
    $F(x_{n})-F(\Phi(x_{n}))=\sum_{j=1}^{r}(f(x_{j,n})-f(\varphi(x_{j,n})))\qquad\forall n\in\mathbb{N}$
    and since each of the summands in the above sum is non-negative, it follows that
    $\lim_{n\to\infty}\left(f(x_{j,n})-f(\varphi(x_{j,n}))\right)=0\qquad\forall j\in\{1,2,\dots,r\}.$
    Moreover, since $(F(x_{n}))_{n}$ is bounded and since $f$ is bounded from below, it follows that $(f(x_{j,n}))_{n}$ is bounded for each $j$. By assumption, $(\varphi,f)$ satisfies condition ($D_{2}$). In particular, it follows that $(x_{1,n})_{n}$ has a convergent subsequence $(x_{1,n_{k}})_{n\in\mathbb{N}}$. Likewise, the sequence $(x_{2,n_{k}})_{k}$ has a convergent subsequence $(x_{2,n_{k_{\ell}}})_{\ell}$. Going on like this and passing to further subsequences until we reach the $r$-th component sequence, we eventually obtain a subsequence of $(x_{n})_{n}$, which we simply denote by $(y_{m})_{m\in\mathbb{N}}$ with the property that all of its component sequences converge. Thus, $(y_{m})_{m\in\mathbb{N}}$ is a convergent subsequence of $(x_{n})_{n\in\mathbb{N}}$ whose limit lies in $\overline{B}\cap\mathrm{Fix\;}\Phi_{1}$ by property ($D_{1}$). This shows that $(\Phi,F)$ has property ($D_{2}$).
  2. b)
    By assumption the restriction $\Phi|_{A}\colon A\to A$ is well-defined. Since $(\Phi,F)$ satisfies condition ($D_{1}$) by a), it is evident that $(\Phi|_{A},F|_{A})$ satisfies ($D_{1}$).
    Let $B\subset A$ be a subset, such that $F$ is bounded on $B$ and
    $0=\inf\{F|_{A}(x)-F|_{A}(\Phi(x))\mid x\in B\}=\inf\{F(x)-F(\Phi(x))\mid x\in B\}.$
    Since $(\Phi,F)$ satisfies ($D_{2}$), it holds that $\overline{B}\cap\mathrm{Fix\;}\Phi\neq\emptyset$. We derive from (4.1) that
    $\overline{B}\cap\mathrm{Fix\;}\Phi\subset\overline{A}\cap\mathrm{Fix\;}\Phi\subset A,$
    such that $\operatorname{cl}_{A}(B)\cap\mathrm{Fix\;}\Phi=\overline{B}\cap\mathrm{Fix\;}\Phi\neq\emptyset$, where $\operatorname{cl}_{A}(B)$ denotes the closure of $B$ in $A$ and $\overline{B}$ the closure of $B$ in $X$. It follows from this observation that $(\Phi|_{A},F|_{A})$ satisfies $(D_{2})$.
Theorem 4.2.
Let $X$ be a path-connected ANR, let $\varphi\colon X\to X$ be continuous and homotopic to the identity, let $f\colon X\to\mathbb{R}$ be a Lyapunov function for $\varphi$ which is bounded from below and let $\lambda\in\mathbb{R}$. Assume that $f^{\lambda}$ is path-connected, that $(\varphi,f)$ satisfies condition (D) and that $f(\mathrm{Fix\;}\varphi)$ is discrete. For all $\lambda\in\mathbb{R}$ it holds that
$\operatorname{\mathsf{TC}}_{r}(f^{\lambda})\leq\sum_{\mu\in(-\infty,r\lambda]}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,f^{\lambda}}\Big{(}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\varphi)\Big{)}\ \Big{|}\ \mu_{1},\dots,\mu_{r}\in(-\infty,\lambda],\,\sum_{i=1}^{r}\mu_{i}=\mu\Big{\}}$
Proof.
Choose $\lambda\in\mathbb{R}$ and put $Y:=f^{\lambda}\subset X$. Since $f$ is a Lyapunov function for $\varphi$, it holds that $\varphi(Y)\subset Y$. Let again $\Phi:=\varphi^{r}\colon X^{r}\to X^{r}$, such that $F$ is a Lyapunov function for $\Phi$. Then $\Phi(Y^{r})\subset Y^{r}$ and since $Y^{r}$ is closed, we derive from Lemma 4.1.b) that $(\Phi|_{Y^{r}},F|_{Y^{r}})$ satisfies condition (D). By definition of $F$, it holds that $Y^{r}\subset F^{r\lambda}$. In terms of the restriction $F|_{Y^{r}}\colon Y^{r}\to\mathbb{R}$, this yields that $Y^{r}=(F|_{Y^{r}})^{r\lambda}$, so we derive from Theorem 2.10.a) that
$\operatorname{\mathsf{TC}}_{r}(Y)$ $=\operatorname{\mathsf{TC}}_{r,Y}(Y^{r})=\operatorname{\mathsf{TC}}_{r,Y}((F|_{Y^{r}})^{r\lambda})\leq\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,Y}((F|_{Y^{r}})_{\mu}\cap\mathrm{Fix\;}(\Phi|_{Y^{r}}))$
$=\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,Y}(F_{\mu}\cap Y^{r}\cap\mathrm{Fix\;}\Phi).$
Apparently, since $f$ is bounded from below, it holds for each $\mu\in\mathbb{R}$ that
$F_{\mu}\cap Y^{r}$ =$=\{(x_{1},\dots,x_{r})\in X^{r}\mid f(x_{1})+f(x_{2})+\dots+f(x_{r})=\mu\}\cap Y^{r}$
$=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,r\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap Y)=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}f_{\mu_{i}}$
Since $\mathrm{Fix\;}\Phi=(\mathrm{Fix\;}\varphi)^{r}$, we derive that
$F_{\mu}\cap Y^{r}\cap\mathrm{Fix\;}\Phi=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\varphi).$
By assumption on $f$, the latter is a union of finitely many pairwise disjoint closed subsets of $X^{r}$. Thus, iterating Proposition 2.6.d) in the above computation yields
$\operatorname{\mathsf{TC}}_{r}(Y)$ $\leq\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,Y}\Big{(}\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\varphi)\Big{)}$
$=\sum_{\mu\in(-\infty,r\lambda]}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,Y}\Big{(}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\varphi\Big{)}\ \Big{|}\ \mu_{1},\dots,\mu_{r}\in(-\infty,\lambda],\ \sum_{i=1}^{r}\mu_{i}=\mu\Big{\}},$
which we wanted to show. ∎
Corollary 4.3.
Let $(X,g)$ be a complete Riemannian Banach manifold, let $f\in C^{1}(X)$ be bounded from below and let $\lambda\in\mathbb{R}$. If $(f,g)$ satisfies the Palais-Smale condition and if $f(\operatorname{Crit}f)$ is discrete, then
$\operatorname{\mathsf{TC}}_{r}(f^{\lambda})\leq\sum_{\mu\in(-\infty,r\lambda]}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,f^{\lambda}}(\operatorname{Crit}_{\mu_{1}}f\times\dots\times\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \mu_{1},\dots,\mu_{r}\in(-\infty,\lambda],\,\sum_{i=1}^{r}\mu_{i}=\mu\Big{\}}.$
Proof.
This follows from Theorem 4.2 in the same way that Corollary 2.11.a) follows from Theorem 2.10. ∎
5. Sequential parametrized subspace complexities as index functions
In the same sense that Rudyak generalized the notion of topological complexity to sequential topological complexities, the notion of parametrized topological complexity as introduced by Cohen, Farber and Weinberger in [CFW21] was generalized to sequential parametrized topological complexities by Farber and Paul in [FP22]. It has since been further studied by the same authors in [FP23] and by Farber and Oprea in [FO24].
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$ and let $p\colon E\to X$ be a fibration with path-connected fibers. We further let $E^{I}_{X}$, $E^{r}_{X}$ and $\Pi_{r}$ be defined aas in the introduction.
In analogy with the line of argument of section 2, we want to construct an index function on $E^{r}_{X}$ out of subspace versions of sequential parametrized TCs.
Definition 5.1.
  1. a)
    Let $A\subset E^{r}_{X}$. A local section $s\colon A\to E^{I}_{X}$ of $\Pi_{r}$ is called a sequential parametrized motion planner on $A$
  2. b)
    The $r$-th sequential parametrized subspace complexity of $A$, denoted by $\operatorname{\mathsf{TC}}_{r,p}(A)$, is given as the minimal value of $n\in\mathbb{N}$ for which there exist open subsets $U_{1},U_{2},\dots,U_{n}\subset E^{r}_{X}$ with $A\subset\bigcup_{i=1}^{n}U_{i}$, such that there exists a continuous sequential parametrized motion planner $s_{i}\colon U_{i}\to PX$ for each $i\in\{1,2,\dots,n\}$. In case there is no such $n$, we put $\operatorname{\mathsf{TC}}_{r,X}(A):=\infty$
Throughout the following we put
$q_{r}\colon E_{X}^{r}\to X,\qquad q_{r}(u_{1},\dots,u_{r}):=p(u_{1}).$
More abstractly, $q_{r}$ is given as the pullback fibration
$q_{r}=\Delta_{r}^{*}p^{r},$
where $\Delta_{r}\colon X\to X^{r}$, $\Delta(x)=(x,x,\dots,x)$ and where $p^{r}\colon E^{r}\to X^{r}$ denotes the $r$-fold cartesian product of $p$ with itself. In particular, $q_{r}$ is itself a fibration.
The following lemma is a straightforward analogue of Lemma 2.2 with respect to parametrized motion planning.
Lemma 5.2.
Assume that $E$ and $X$ are Hausdorff spaces and let $d_{E}\colon E^{r}_{X}\to E^{r}_{X}$, $d_{E}(u_{1},u_{2},\dots,u_{r})=(u_{1},u_{1},\dots,u_{1})$. Let $A\subset E^{r}_{X}$ be a subset. The following statements are equivalent:
  1. (1)
    $A$ admits a continuous sequential parametrized motion planner.
  2. (2)
    There exists a deformation $h\colon A\times[0,1]\to E^{r}_{X}$ with $h(a,1)=d_{E}(a)$ for each $a\in A$, such that
    $q_{r}(h(a,t))=q_{r}(a)\qquad\forall a\in A,\ t\in[0,1].$
Proof.
(1) $\Rightarrow$ (2): One constructs a deformation by the same formula as in the proof of the implication “(1) $\Rightarrow$ (2)” of Lemma 2.2 and checks that it has the desired properties.
(2) $\Rightarrow$ (1):  Let $h\colon A\times[0,1]\to E^{r}_{X}$ be a fiber-preserving deformation with respect to $q_{r}$. We further let $h_{0},\dots,h_{r-1}\colon A\times[0,1]\to E$ and $s\colon A\to PE$ be defined as in (2.1). Since $E$ is Hausdorff, $E^{r}_{X}$ is a Hausdorff space, hence $s$ is continuous. Since $h$ is fiber-preserving with respect to $q_{r}$, it holds that $p(h_{j}(a,t))=p(h_{j}(a,0))=p(h_{0}(a,0))$ for all $a\in A$, $t\in[0,1]$ and $j\in\{0,1,\dots,r-1\}$. Consequently
$p((s(a))(t))=p(h_{0}(a,0))\quad\forall a\in A,\ t\in[0,1],$
such that $s(a)\in E^{I}_{X}$ for all $a\in A$. Thus, $s$ is indeed a continuous sequential parametrized motion planner. ∎
In section 2 we have restricted ourselves to considering ANRs when considering sequential subspace complexities, as we could establish the continuity property of an index function under this assumption. Likewise, we need to make a similar assumption on our fibration $p\colon E\to X$ to ensure that sequential parametrized subspace complexities will have the continuity property. For this purpose, we next introduce an analogue to the definition of an ANR for fibrations that was originally defined by Yagasaki in [Yag85] and that can be seen as an ANR-type property for fiberwise spaces.
Definition 5.3.
Let $Q$ and $B$ be topological spaces and let $f\colon Q\to B$ be continuous. We call $f$ an absolute neighborhood fiber retract (ANFR) if $Q$ and $B$ are metrizable and if $f$ satisfies the following condition:
Let $T$ be a metrizable space, let $q\colon T\to B$ be continuous and let $A\subset T$ be closed. Then for any continuous map $g\colon A\to Q$ with $q|_{A}=f\circ g$ there are an open neighborhood $U\subset T$ of $A$ and a continuous map $G\colon U\to Q$ with $G|_{A}=g$ and $q|_{U}=f\circ G$.
Remark 5.4.
The notion of an ANFR from Definition 5.3 slightly differs from the one introduced in [Yag85]. In his terminology, our definition is the one of an absolute neighborhood fiber extensor. However, Yagasaki has shown in [Yag85, Proposition 1.1] that both notions are indeed equivalent.
The following result yields two helpful criteria to show the ANFR property for fibrations.
Lemma 5.5 ([Yag85, 1.2.(iii) and (v)]).
Let $Q$ and $B$ be metrizable spaces and let $f\colon Q\to B$ be a fibration.
  1. a)
    If $Q$ and $B$ are ANRs, then $f$ is an ANFR.
  2. b)
    If $f$ is a fiber bundle with typical fiber $F$ and if $F$ is an ANR, then $f\colon Q\to B$ is an ANFR.
Our concrete aim is to show that $q_{r}\colon E^{r}_{X}\to X$ is an ANFR for which we need the following auxiliary lemma.
Lemma 5.6.
Let $Q$ and $B$ be ANRs and let $f\colon Q\to B$ be a fibration. If $C\subset B$ is an ANR, then $f^{-1}(C)$ is an ANR.
Proof.
Put $D:=f^{-1}(C)$. Let $T$ be a metrizable space, let $A\subset T$ be closed and let $g\colon A\to D$ be continuous. Then $f\circ g\colon A\to B$ is continuous with $(f\circ g)(A)\subset C$. Since $C$ is an ANR, there exist an open neighborhood $V\subset T$ of $A$ and a continuous map $q\colon V\to B$ with $q|_{A}=f\circ g$ and $q(V)\subset C$. As an open subset of $T$, $V$ is again metrizable. Since $Q$ and $B$ are ANRs, $p$ is an ANFR by Lemma 5.5.a). Thus, there exist a neighborhood $U\subset V$ of $A$, which is open in $T$ and a continuous map $G\colon U\to Q$, such that
$G|_{A}=g\qquad\text{and}\qquad q|_{U}=f\circ G.$
In particular, it holds that $(f\circ G)(U)\subset q(U)\subset C$, which shows that $G(U)\subset D$. So we can restrict the codomain of $G$ to $G\colon U\to D$, which yields the desired continuous extension of $g$. Hence, $D$ is an ANR. ∎
Proposition 5.7.
If $E$ and $X$ are ANRs, then $E^{I}_{X}$ and $E^{r}_{X}$ are ANRs.
Proof.
Since $E$ and $X$ are ANRs, $E^{r}$ and $X^{r}$ are again ANRs. Moreover, $p^{r}\colon E^{r}\to X^{r}$ is a fibration for which $E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$. Since $\Delta_{r,X}$ is homeomorphic to $X$, it is an ANR as well. As a special case of Lemma 5.6, we thus obtain that $E^{r}_{X}$ is an ANR.
Considering $E^{I}_{X}$, we first observe that since $p$ is a fibration the induced map of path spaces
$p_{\#}\colon PE\to PX,\quad p_{\#}(\gamma)=p\circ\gamma,$
is a fibration as well. Since $E$ and $X$ are ANRs, it further follows from [Hu65, Theorem VI.2.4] that $PE$ and $PX$ are ANRs. By definition of $E^{I}_{X}$, it holds that
$E^{I}_{X}=p_{\#}^{-1}(c(X)),$
where $c(X)\subset PX$ denotes the space of constant paths in $X$. Since $c(X)$ is homeomorphic to $X$, it is an ANR, so we again obtain from Lemma 5.6 that $E^{I}_{X}$ is an ANR. ∎
The following lemma is the analogue of Lemma 2.5 for the parametrized setting.
Lemma 5.8.
Assume that $E$ and $X$ are ANRs and let $A\subset E^{r}_{X}$ be a closed subset which admits a continuous sequential parametrized motion planner. Then there exists an open neighborhood $U\subset E^{r}_{X}$ of $A$ which admits a continuous sequential parametrized motion planner.
Proof.
By Proposition 5.7, $E^{r}_{X}$ is an ANR, so it follows from Lemma 5.5.a) that $q_{r}\colon E^{r}_{X}\to X$ is an ANFR. By Lemma 5.2, there exists a deformation $h^{\prime}\colon A\times[0,1]\to E^{r}_{X}$ with $h^{\prime}(a,1)=d_{E}(a)$ and $q_{r}(h^{\prime}(a,t))=q_{r}(a)$ for all $a\in A$ and $t\in[0,1]$, where $d_{E}(u_{1},u_{2},\dots,u_{r})=(u_{1},u_{1},\dots,u_{1})$. We put $Q^{\prime}:=(E^{r}_{X}\times\{0,1\})\cup(A\times[0,1])$ and let $H^{\prime}\colon Q^{\prime}\times[0,1]\to E^{r}_{X}$ be defined as in (2.2), only replacing $d$ by $d_{E}$ in the second case. Then $q_{r}\circ\mathrm{pr}_{1}|_{Q}=q_{r}\circ H^{\prime}$, where $\mathrm{pr}_{1}\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ denotes the projection onto the first factor. Since $q_{r}$ is an ANFR, there exist an open neighborhood $V^{\prime}\subset E_{X}^{r}\times[0,1]$ of $Q^{\prime}$ and a continuous $K^{\prime}\colon V^{\prime}\to E_{X}^{r}$, such that $K^{\prime}|_{Q^{\prime}}=H$ and $q_{r}\circ\mathrm{pr}_{1}|_{V^{\prime}}=q_{r}\circ K^{\prime}$.
Arguing as in the proof of Lemma 2.5, we obtain an open subset $U^{\prime}\subset E^{r}_{X}$ with $U^{\prime}\times[0,1]\subset V^{\prime}$ and $A\subset U^{\prime}$. Then $K^{\prime}|_{U^{\prime}\times[0,1]}$ is a deformation with $K^{\prime}(u,1)=d_{E}(u)$ and $(q_{r}\circ K^{\prime})(u,t)=q_{r}(u)$ for all $u\in U^{\prime}$ and $t\in[0,1]$. The claim then follows from Lemma 5.2. ∎
Proposition 5.9.
  1. a)
    For all $A\subset B\subset E^{r}_{X}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,p}(A)\leq\operatorname{\mathsf{TC}}_{r,p}(B).$
  2. b)
    For all $A,B\subset E^{r}_{X}$ it holds that
    $\operatorname{\mathsf{TC}}_{r,p}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,p}(A)+\operatorname{\mathsf{TC}}_{r,p}(B).$
  3. c)
    Let $A\subset E^{r}_{X}$ be a closed subset. Then there exists an open subset $U\subset E^{r}_{X}$ with $A\subset U$ and
    $\operatorname{\mathsf{TC}}_{r,p}(A)=\operatorname{\mathsf{TC}}_{r,p}(U).$
  4. d)
    Assume that $E$ is normal. Let $A,B\subset E^{r}_{X}$ be closed with $A\cap B=\emptyset$. Then
    $\operatorname{\mathsf{TC}}_{r,p}(A\cup B)=\max\{\operatorname{\mathsf{TC}}_{r,p}(A),\operatorname{\mathsf{TC}}_{r,p}(B)\}.$
Proof.
All parts of the proposition follow in complete analogy with the corresponding parts of Proposition 2.6, where we use that by assumption, the fibers of $p$ are path-connected. Here, in the proof of d) we use that since $E$ is normal, $E^{r}$ is normal as well and so is its closed subspace $E^{r}_{X}$. ∎
Proposition 5.10.
Assume that $E$ and $X$ are ANRs, let $A\subset E^{r}_{X}$ be closed and let $\Phi\colon A\times[0,1]\to E^{r}_{X}$ be a deformation of $A$ for which $\Phi_{1}(A)$ is closed and which satisfies
(5.1) $q_{r}(\Phi(a,t))=q_{r}(a)\quad\forall a\in A,\ t\in[0,1].$
Then
$\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))\geq\operatorname{\mathsf{TC}}_{r,p}(A).$
Proof.
Put $k:=\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))$. In analogy with the proof of Proposition 2.7, one shows that there are closed subsets $B_{1},B_{2},\dots,B_{k}\subset\Phi_{1}(A)$ with $\bigcup_{j=1}^{k}B_{j}=\Phi_{1}(A)$, such that for each $j\in\{1,\ldots,k\}$ there exists a sequential parametrized motion planner $s_{j}\colon B_{j}\to PE$. Put $B_{j}^{\prime}:=\Phi_{1}^{-1}(B_{j})$ for each $j\in\{1,\ldots,k\}$, such that $A\subset\bigcup_{j=1}^{k}B_{j}^{\prime}$, and define a map $s_{j}^{\prime}\colon B_{j}^{\prime}\to PE$ by the same expression as in (2.3) in the proof of Proposition 2.7. One checks that condition (5.1) ensures that each $s_{j}^{\prime}$ is indeed a continuous sequential parametrized motion planner. Since $E$ and $X$ are ANRs, it follows from Lemma 5.8 that there are open subsets $U_{1},U_{2}\dots,U_{k}\subset E^{r}_{X}$ with $B^{\prime}_{j}\subset U_{j}$ for each $j\in\{1,2,\dots,k\}$, such that each $U_{j}$ admits a continuous sequential parametrized motion planner. Since $A\subset\bigcup_{j=1}^{k}U_{j}$, this shows that $\operatorname{\mathsf{TC}}_{r,p}(A)\leq k$. ∎
Corollary 5.11.
Assume that $E$ and $X$ are ANRs and let $\Phi\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ be continuous with
$\Phi(u,0)=u,\qquad q_{r}(\Phi(u,t))=q_{r}(u)\qquad\forall u\in E^{r}_{X},\ t\in[0,1].$
Then $\operatorname{\mathsf{TC}}_{r,p}\colon\mathfrak{P}(E^{r}_{X})\to\mathbb{N}\cup\{\infty\}$ is a $\Phi_{1}$-supervariant index function on $E^{r}_{X}$, where $\Phi_{1}(u):=\Phi(u,1)$.
Proof.
It follows immediately from Proposition 5.9.a)-c) that $\operatorname{\mathsf{TC}}_{r,p}$ is an index function. If $A\subset E^{r}_{X}$ is closed, then $\Phi|_{A\times[0,1]}$ is a deformation of $A$ and it follows from Proposition 5.10 that $\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))\geq\operatorname{\mathsf{TC}}_{r,p}(A)$. Thus, $\operatorname{\mathsf{TC}}_{r,p}$ is $\Phi_{1}$-supervariant. ∎
Using Corollary 5.11 we obtain the following result as a direct application of Theorem 1.4.
Theorem 5.12.
Let $E$ and $X$ be ANRs and let $\Phi\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ be continuous with
(5.2) $\Phi(u,0)=u,\quad q_{r}(\Phi(u,t))=q_{r}(u)\quad\forall u\in E^{r}_{X},\ t\in[0,1].$
Let $F\colon E^{r}_{X}\to\mathbb{R}$ be a Lyapunov function for $\Phi_{1}:=\Phi(\cdot,1)$ which is bounded from below and such that $(\Phi_{1},F)$ satisfies condition (D). Then
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\mathrm{Fix\;}\Phi_{1})\qquad\forall\lambda\in\mathbb{R}.$
6. Vertically proportional vector fields and parametrized subspace complexities
We want to study how Theorem 5.12 can be applied to critical points of differentiable functions on $E^{r}_{X}$, assuming that the latter is a smooth manifold. Note that the theorem is not directly applicable to flows of negative pseudo-gradients as such a flow will in general not be fiber-preserving in the sense of condition (5.2). We will overcome this problem by working with flows of vertical components of gradients on the one hand and with a strongly restricted class of functions on the other hand.
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$, let $E$ and $X$ be paracompact Banach manifolds and let $p\colon E\to X$ be a smooth fibration and a surjective submersion whose fibers are path-connected.
Note that $E$ and $X$ are path-connected ANRs by [Pal66a, Cor. to Theorem 5]. Since $p^{r}\colon E^{r}\to X^{r}$ is submersive as well, it follows that
$E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$
is a smooth Banach manifold whose tangent space at $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$ is given by
(6.1) $T_{u}E^{r}_{X}=\{(v_{1},v_{2}\dots,v_{r})\in T_{u}E^{r}\mid Dp_{u_{1}}(v_{1})=Dp_{u_{2}}(v_{2})=\dots=Dp_{u_{r}}(v_{r})\}.$
We recall that given a smooth fibration $f\colon M\to N$ the vertical tangent space of $f$ at a point $m\in M$ is given by
$\mathrm{Ver}_{m}(f):=\ker Df_{m}.$
In particular, $\mathrm{Ver}_{m}(f)$ is always a closed linear subspace of $T_{m}M$. Evidently, it holds for each $u=(u_{1},u_{2},\dots,u_{r})\in E^{r}_{X}$ that
$\mathrm{Ver}_{u}(p^{r})=\ker(Dp^{r})_{u}=\prod_{i=1}^{r}\ker Dp_{u_{i}}=\prod_{i=1}^{r}\mathrm{Ver}_{u_{i}}(p).$
Let again $q_{r}\colon E^{r}_{X}\to X$, $q_{r}(u_{1},\dots,u_{r})=p(u_{1})$, which is a smooth fibration and a surjective submersion as well. Moreover, for each $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$ it holds that
$\mathrm{Ver}_{u}(q_{r})$ =$=\{v\in T_{u}E^{r}_{X}\mid(Dq_{r})_{u}(v)=0\}$
=$=\{(v_{1},\dots,v_{r})\in T_{u}E^{r}\mid 0=Dp_{u_{1}}(v_{1})=Dp_{u_{2}}(v_{2})=\dots=Dp_{u_{r}}(v_{r})\}$
$=\mathrm{Ver}_{u}(p^{r}).$
Thus, we have shown that
(6.2) $\mathrm{Ver}_{u}(q_{r})=\prod_{i=1}^{r}\mathrm{Ver}_{u_{i}}(p)\qquad\forall u=(u_{1},u_{2},\dots,u_{r})\in E^{r}_{X}.$
The following result is an immediate consequence of the results of the previous section. We recall that a vector field $Y\colon M\to TM$ on a smooth manifold $M$ is called forward-complete if it admits a flow on all of $M\times[0,+\infty)$.
Theorem 6.1.
Let $V\colon E^{r}_{X}\to TE^{r}_{X}$ be a forward-complete continuous vector field with
$V(u)\in\mathrm{Ver}_{u}(p^{r})\qquad\forall u\in E^{r}_{X}.$
Assume that the time-1 map $\Phi_{1}$ of its flow admits a Lyapunov function $F\colon E^{r}_{X}\to\mathbb{R}$, which is bounded from below, such that $(\Phi_{1},F)$ satisfies condition (D). Then
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\mathrm{sing}(V))\qquad\forall\lambda\in\mathbb{R}.$
Proof.
Let $\Phi\colon E^{r}_{X}\times[0,+\infty)\to E^{r}_{X}$ be the flow of $V$. By assumption and the above computations, it holds that $V(u)\in\mathrm{Ver}_{u}(q_{r})$ for all $u\in E^{r}_{X}$ for each $u\in E^{r}_{X}$. By elementary results for flows of vector fields, this yields
$q_{r}(\Phi(u,t))=q_{r}(u)\quad\forall u\in E^{r}_{X},\ t\in[0,+\infty).$
Thus, since $\Phi_{1}(u)=\Phi(u,1)$ for each $u\in E^{r}_{X}$, we derive from Theorem 5.12 that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\mathrm{Fix\;}\Phi_{1}).$
Since by construction $\mathrm{Fix\;}\Phi_{1}=\mathrm{sing}(V)$, the claim immediately follows. ∎
Definition 6.2.
Let $M$ and $N$ be Banach manifolds, let $q\colon M\to N$ be smooth and let $g$ be a Riemannian metric on $M$.
  1. a)
    We define
    $\mathrm{pr}_{\mathrm{ver}}\colon TM\to TM$
    as the bundle map for which $\mathrm{pr}_{\mathrm{ver}}|_{T_{m}M}$ is the orthogonal projection onto $\mathrm{Ver}_{m}(q)$ with respect to $g$ for each $m\in M$.
  2. b)
    Given a vector field $Y\colon M\to TM$, we define another vector field by
    $Y^{\mathrm{ver}}\colon M\to TM,\quad Y^{\mathrm{ver}}:=\mathrm{pr}_{\mathrm{ver}}\circ Y.$
    Given $F\in C^{1}(M)$, we further put
    $\nabla^{\mathrm{ver}}F:=(\nabla^{g}F)^{\mathrm{ver}},$
    where $\nabla^{g}F\colon M\to TM$ denotes the gradient of $F$ with respect to $g$.
  3. c)
    We call a vector field $Y\colon M\to TM$ vertically proportional with respect to $q$ and $g$ if there exists $C>0$, such that
    $\|Y(x)\|_{g}\leq C\cdot\|Y^{\mathrm{ver}}(x)\|_{g}\qquad\forall x\in M,$
    where $\|\cdot\|_{g}$ denotes the norm induced by $g$ on each $T_{x}M$.
Remark 6.3.
Let $M$, $N$, $g$ and $q$ be given as in Definition 6.2 and let $Y\colon M\to TM$ be a vector field. Apparently, $\mathrm{sing}(Y)\subset\mathrm{sing}(Y^{\mathrm{ver}})$. Conversely, if $Y$ is vertically proportional, then every singular point of $Y^{\mathrm{ver}}$ is a singular point of $Y$ as well. Thus, if $Y$ is vertically proportional, then it necessarily holds that
(6.3) $\mathrm{sing}(Y)=\mathrm{sing}(Y^{\mathrm{ver}}).$
As we will see in the next results, if a function $F\in C^{1}(E^{r}_{X})$ has a vertically proportional gradient with respect to a complete Riemannian metric, then this will allow us to apply Theorem 6.1 to estimate the number of critical points of $F$ by sequential parametrized subspace complexities. However, as condition (6.3) already indicates, the vertical proportionality of its gradient is a strong restriction on the function itself.
Proposition 6.4.
Let $(M,g)$ be a complete Riemannian Banach manifold, let $N$ be a smooth Banach manifold and let $q\colon M\to N$ be smooth. Let $F\in C^{1}(M)$, let $\rho\colon\mathbb{R}\to\mathbb{R}$ and $h\colon M\to\mathbb{R}$ be given as in Example 1.7 and put
$Y\colon M\to TM,\quad Y(m):=h(m)\cdot\nabla^{g}F(m).$
Assume $\nabla^{g}F$ is vertically proportional with respect to $q$ and $g$.
  1. a)
    Then $Y^{\mathrm{ver}}\colon M\to TM$ is a pseudo-gradient vector field for $F$.
  2. b)
    Let $\phi^{\mathrm{ver}}_{1}\colon M\to M$ be the time-1 map of the flow of $-Y^{\mathrm{ver}}$. Then $(\phi^{\mathrm{ver}}_{1},F)$ satisfies condition (D).
Proof.
[…]
$\|Y(m)\|\leq C\cdot\|Y^{\mathrm{ver}}(m)\|\qquad\forall m\in M.$
  1. a)
    Let us check the three properties of a pseudo-gradient for $Y^{\mathrm{ver}}$ as listed in Definition 1.5.
    1. (i)
      Since $\mathrm{pr}_{\mathrm{ver}}$ is fiberwise given by the bounded-linear projection onto $\mathrm{Ver}_{u}(q)$, it follows from the construction of the Sasaki metric that $\mathrm{pr}_{\mathrm{ver}}$ is $1$-Lipschitz-continuous. Thus, it follows from the Lipschitz continuity of $Y$ that $Y^{\mathrm{ver}}$ is Lipschitz-continuous.
    2. (ii)
      Evidently, $\|Y^{\mathrm{ver}}(m)\|\leq\|Y(m)\|$ for each $m\in M$, so the upper bound on $\|Y^{\mathrm{ver}}(m)\|$ follows immediately from the corresponding upper bound on $\|Y\|$, which exists since $Y$ is a pseudo-gradient.
    3. (iii)
      One easily sees that
      $Y^{\mathrm{ver}}(m)=h(m)\cdot\nabla^{\mathrm{ver}}F(m)\qquad\forall m\in M.$
      Thus, we compute for each $m\in M$ that
      $DF_{m}(Y^{\mathrm{ver}}(m))$ $=h(m)\cdot DF_{m}(\nabla^{\mathrm{ver}}F(m))=h(m)\cdot g_{M}(\nabla F(m),\nabla^{\mathrm{ver}}F(m))$
      $=h(m)\cdot g_{M}(\nabla^{\mathrm{ver}}F(m),\nabla^{\mathrm{ver}}F(m))=h(m)\cdot\|\nabla^{\mathrm{ver}}F(m)\|^{2}$
      $\geq h(m)\cdot C^{-2}\|\nabla^{g}F(m)\|^{2}.$
      If $\|\nabla^{g}F(m)\|\leq 1$, then it follows immediately that
      $DF_{m}(Y^{\mathrm{ver}}(m))\geq C^{-2}\min\{\|\nabla^{g}F(m)\|,\|\nabla^{g}F(m)\|^{2}\}.$
      For all $m\in M$ with $\|\nabla^{g}F(m)\|>1$ it holds that $h(m)\geq\frac{1}{\|\nabla^{g}F(m)\|}$ and thus
      $DF_{m}(Y^{\mathrm{ver}}(m))\geq C^{-2}\|\nabla^{g}F(m)\|=C^{-2}\min\{\|\nabla^{g}F(m)\|,\|\nabla^{g}F(m)\|^{2}\}.$
      This shows property (iii) of a pseudo-gradient if we put $C_{2}:=C^{-2}$ in the situation of Definition 1.5.
  2. b)
    This follows immediately from combining part a) with Theorem 1.10.
Throughout the rest of this section, we choose and fix a complete Riemannian metric $g$ on $E$ and let $\bar{g}$ be the Riemannian metric on $E^{r}_{X}$ that is obtained by restricting the product metric $g^{r}$ on $E^{r}$.
Using Proposition 6.4, we can now apply Theorem 6.1 to critical points of functions and thereby establish an analogue to Corollary 2.11 for the parametrized setting.
Theorem 6.5.
Let $F\in C^{1}(E^{r}_{X})$ be bounded from below. Assume that $(F,\bar{g})$ satisfies the PS condition and that $\nabla^{\bar{g}}F$ is vertically proportional with respect to $q_{r}\colon E^{r}_{X}\to X$ and $\bar{g}$. For each $\lambda\in\mathbb{R}$ it holds that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\left\{\operatorname{\mathsf{TC}}_{r,p}(C)\mid C\in\mathcal{C}_{\mu}(F)\right\}.$
Proof.
Let $Y\colon E^{r}_{X}\to TE^{r}_{X}$ be the pseudo-gradient of $F$ that is constructed from $\nabla^{\bar{g}}F$ as in Proposition 6.4. By definition of $Y^{\mathrm{ver}}$, it holds that
[…]
By Proposition 6.4, we can apply Theorem 5.12 to $Y^{\mathrm{ver}}$ with respect to $q_{r}$ and thus obtain for each $\lambda\in\mathbb{R}$ that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\operatorname{Crit}F).$
Using Proposition 5.9.d), the claim is then derived in the same way as Corollary 2.11. ∎
In the rest of this section, we will apply this result to parametrized analogues of navigation functions, which are defined as follows. .
Definition 6.6.
We call a continuously differentiable function $F\colon E^{r}_{X}\to\mathbb{R}$ a parametrized $r$-navigation function if
  1. (i)
    $F$ satisfies the PS condition with respect to a complete Riemannian metric on $E^{r}_{X}$,
  2. (ii)
    $F(u)\geq 0$ for all $u\in E^{r}_{X}$ and $F^{-1}(\{0\})=\Delta_{r,E}$.
Remark 6.7.
  1. (1)
    Evidently, if $F\colon E^{r}\to\mathbb{R}$ is an $r$-navigation function in the sense of Definition 3.1, then $F|_{E^{r}_{X}}$ will be a parametrized $r$-navigation function.
  2. (2)
    In analogy with Remark 3.2.(2), one argues that if $E$ and $X$ are closed finite-dimensional manifolds, then any continuously differentiable $F\colon E^{r}_{X}\to\mathbb{R}$ which has property (ii) of Definition 6.6 will be a parametrized $r$-navigation function.
Without additional difficulties, one can now establish an analogue of Theorem 3.3.
Theorem 6.8.
Let $F\in C^{1}(E^{r}_{X})$ be a parametrized $r$-navigation function for which $\nabla^{\bar{g}}F$ is vertically proportional with respect to $q_{r}$ and $\bar{g}$. Then
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}_{r,p}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1\qquad\forall\lambda\in\mathbb{R}.$
Proof.
This follows from Theorem 6.5 in the same way as Theorem 3.3 is derived from Corollary 2.11, since the sequential motion planner by constant paths on $\Delta_{r,E}$ is indeed a continuous sequential parametrized motion planner which extends to an open neighborhood of $\Delta_{r,E}$ by Lemma 5.8. ∎
In the following, we present a method of obtaining parametrized navigation functions which satisfy the requirements of Theorem 6.8.
Let $E$ and $B$ be closed smooth manifolds and let $p\colon E\to B$ be a smooth fibration. Since a smooth fibration between finite-dimensional manifolds is necessarily a submersion, $E\times_{B}E$ is a smooth manifold as well. For suitable numbers $N_{1},N_{2}\in\mathbb{N}$ we choose smooth embeddings $i_{B}\colon B\hookrightarrow\mathbb{R}^{N_{1}}$ and $i_{E}\colon E\hookrightarrow\mathbb{R}^{N_{2}}$. Then one easily checks that
$i_{0}\colon E\times_{B}E\to\mathbb{R}^{N_{1}+2N_{2}},\qquad i_{0}(e_{1},e_{2})=(i_{B}(p(e_{1})),i_{E}(e_{1}),i_{E}(e_{2})),$
is a smooth embedding as well.
Recall that if $\pi\colon M\to N$ is a fibration between smooth manifolds, where $M$ is equipped with a Riemannian metric we define the horizontal subspace at $x\in M$ by
$\mathrm{Hor}_{x}(\pi):=\big{(}\mathrm{Ver}_{x}(q)\big{)}^{\perp}\subset T_{x}M.$
Proposition 6.9.
Consider the function
$F\colon E\times_{B}E\to\mathbb{R},\qquad F(e_{1},e_{2})=\|i_{E}(e_{1})-i_{E}(e_{2})\|^{2},$
[…]
  1. a)
    Then $F$ is a parametrized $2$-navigation function.
  2. b)
    Let $g_{0}$ be the Riemannian metric on $E\times_{B}E$ for which $i_{0}$ is an isometric embedding with respect to the Euclidean metric on $\mathbb{R}^{N_{1}+2N_{2}}$. Assume that for all $(e_{1},e_{2})\in E\times_{B}E$ and $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$ we have
    (6.4) $\langle i_{E}(e_{1})-i_{E}(e_{2}),u_{1}-u_{2}\rangle=0$
    where we write $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{2})\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{2n_{2}}$. Then $\nabla^{g_{0}}F$ is vertically proportional with respect to $q\colon E\times_{B}E\to B$, $q(e_{1},e_{2})=p(e_{1})$.
Proof.
  1. a)
    Apparently, $F$ is smooth with $F(e_{1},e_{2})\geq 0$ for all $(e_{1},e_{2})\in E\times_{B}E$ and $F(e_{1},e_{2})=0$ holds if and only if $e_{1}=e_{2}$, i.e. if $(e_{1},e_{2})\in\Delta_{2,E}$. Since $E\times_{B}E$ is compact, this yields that $F$ is a navigation function, see Remark 6.7.(2).
  2. b)
    To compute $\nabla^{g_{0}}F$, we consider the function $\varphi\colon\mathbb{R}^{N_{2}}\times\mathbb{R}^{N_{2}}\to\mathbb{R}$, $\varphi(x,y)=\|x-y\|^{2}$, which is easily seen to be differentiable with
    $D\varphi_{(x,y)}(v,w)=2\left<x-y,v-w\right>\qquad\forall x,y,v,w\in\mathbb{R}^{N_{2}}$
    with respect to the Euclidean inner product. With $\mathrm{pr}_{2}\colon\mathbb{R}^{N_{1}}\times\mathbb{R}^{2N_{2}}\to\mathbb{R}^{2N_{2}}$ denoting the projection, we observe that
    $F=\varphi\circ\mathrm{pr}_{2}\circ i_{0}$
    and thus
    (6.5) $DF_{(e_{1},e_{2})}=D\varphi_{(i_{E}(e_{1}),i_{E}(e_{2}))}\circ(D\mathrm{pr}_{2})_{i_{0}(e_{1},e_{2})}\circ(Di_{0})_{(e_{1},e_{2})}$
    Let $(e_{1},e_{2})\in E\times_{B}E$ and let $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$. We write $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{2})\in\mathbb{R}^{n_{1}+2n_{2}}$. Set $(x_{1},x_{2})=(i_{E}(e_{1}),i_{E}(e_{2}))\in\mathbb{R}^{2n_{2}}$. Then equation (6.5) yields
    $DF_{(e_{1},e_{2})}(u)=D\varphi_{(x_{1},x_{2})}(u_{1},u_{2})=2\langle x_{1}-x_{2},u_{1}-u_{2}\rangle,$
    which vanishes by assumption. It follows that $\nabla^{g_{0}}F(e_{1},e_{2})\in\mathrm{Ver}_{(e_{1},e_{2})}(q)$ for all $(e_{1},e_{2})\in E\times_{B}E$. We immediately derive that $\nabla^{g_{0}}F(e_{1},e_{2})=\nabla^{\mathrm{ver}}F(e_{1},e_{2})$ for all $(e_{1},e_{2})\in E\times_{B}E$. In particular, $\nabla^{g_{0}}F$ is vertically proportional.
Remark 6.10.
Note that condition (6.4) in part b) of Proposition 6.9 is satisfied if for all $(e_{1},e_{2})\in E\times_{B}E$ and $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$ we have that $(Di_{0})_{(e_{1},e_{2})}(u)\in\mathbb{R}^{n_{1}}\times\Delta_{\mathbb{R}^{n_{2}}}$, i.e. it is of the form $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{1})$ for some $u_{0}\in\mathbb{R}^{n_{1}}$ and some $u_{1}\in\mathbb{R}^{n_{2}}$.
7. Sequential parametrized TC of restrictions of fibrations to sublevel sets
We want to apply the reults from the previous section in the case that a smooth fibration $p\colon E\to X$ between Banach manifolds and a function $f\in C^{1}(E)$ are given, such that the restriction $p|_{f^{\lambda}}\colon f^{\lambda}\to X$ is again a fibration. We note that this is a strong assumption, but we will establish a class of fibrations having this property in section 8 below.
Here, we want to derive upper bounds on $\operatorname{\mathsf{TC}}[p|_{f^{\lambda}}\colon f^{\lambda}\to X]$ under the fibration assumption, i.e. upper bounds on sequential parametrized TCs instead of sequential parametrized subspace complexities. Such a result can be seen as a “parametrized analogue” of Corollary 4.3 on sublevel sets.
For this purpose, we will study functions $F\in C^{1}(E^{r}_{X})$ for which there exists $f\in C^{1}(E)$, such that
$F(u_{1},\dots,u_{r})=f(u_{1})+\dots+f(u_{r}).$
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$, let $E$ and $X$ be paracompact Banach manifolds and let $p\colon E\to X$ be a smooth fibration and a surjective submersion whose fibers are path-connected. We further let $g$ be a Riemannian metric on $E$ and let $\bar{g}$ be the Riemannian metric on $E^{r}_{X}$ that is obtained by restricting the product metric $g^{r}$ on $E^{r}$. Given $A_{1},\dots,A_{r}\subset E$, we will further denote
$A_{1}\times_{X}A_{2}\times_{X}\dots\times_{X}A_{r}:=\Big{\{}(a_{1},\dots,a_{r})\in\prod_{i=1}^{r}A_{i}\ \Big{|}\ p(a_{1})=\dots=p(a_{r})\Big{\}}=E^{r}_{X}\cap\prod_{i=1}^{r}A_{i}.$
As a first step, we want to investigate the critical points of such functions as well as the vertical proportionality of their gradients. We first introduce another condition on the critical points of a function in a general setting.
Definition 7.1.
Let $M$ and $N$ be Banach manifolds, let $q\colon M\to N$ be a smooth submersion and let $M_{x}:=q^{-1}(\{x\})$ denote the fiber of $q$ over $x\in N$. We say that a function $F\in C^{1}(M)$ satisfies condition (A) if the following is satisfied:
(A)  Let $x\in N$ and let $u\in M_{x}$ be a critical point of $F|_{M_{x}}$. Then $u$ is a critical point of $F$.
Remark 7.2.
Let $M$ and $N$ be Banach manifolds, $\rho$ be a Riemannian metric on $M$, $q:M\to N$ be a smooth submersion and $F\in C^{1}(M)$. One checks without difficulties that $\nabla^{\mathrm{ver}}F(u)$ coincides with the gradient of $F|_{M_{q(u)}}$ at $u\in M$ with respect to $\rho|_{M_{q(u)}}$. Thus, $F$ satisfies condition (A) if and only if for each $u\in M$ it holds that
$\nabla^{\mathrm{ver}}F(u)=0\quad\Rightarrow\quad\nabla^{\rho}F(u)=0.$
From this observation it is evident that if $\nabla^{\rho}F$ is vertically proportional, then $F$ necessarily satisfies condition (A).
The reason for studying condition (A) in our context is that it allows us to understand the critical points of functions on $E^{r}_{X}$ that arise as restrictions of products in a better way and without studying any Riemannian metric.
Proposition 7.3.
Let $f\in C^{1}(E)$ and consider $F\colon E^{r}_{X}\to\mathbb{R}$, $F(u_{1},\dots,u_{r}):=\sum_{i=1}^{r}f(u_{i})$. If $f$ satisfies condition (A), then
$\operatorname{Crit}F=\operatorname{Crit}f\times_{X}\operatorname{Crit}f\times_{X}\dots\times_{X}\operatorname{Crit}f.$
Proof.
The inclusion“$\supset$” is evident, so it only remains to show that every critical point of $F$ arises in this way. Let $u=(u_{1},\dots,u_{r})\in\operatorname{Crit}F$, such that
$DF_{u}(v_{1},\dots,v_{r})=\sum_{i=1}^{r}Df_{u_{i}}(v_{i})=0\qquad\forall(v_{1},\dots,v_{r})\in T_{u}E^{r}_{X}.$
Let $X_{i}:=p^{-1}(\{p(u_{i})\})$ for each $i\in\{1,2,\dots,r\}$. Since $(0,\dots,0,v_{i},0,\dots,0)\in T_{u}E^{r}_{X}$ for all $v_{i}\in\mathrm{Ver}_{u_{i}}(p)$, it particularly follows for each $i$ that
$Df_{u_{i}}(v_{i})=0\quad\forall v_{i}\in\mathrm{Ver}_{u_{i}}(p)\quad\Rightarrow u_{i}\in\operatorname{Crit}(f|_{X_{i}}).$
Since $f$ satisfies condition (A), this yields that $u_{i}\in\operatorname{Crit}f$ for each $i\in\{1,2,\dots,r\}$, which shows the claim. ∎
Next we want to investigate the vertical proportionality of gradients of such functions with respect to $q_{r}$. For this purpose we establish an auxiliary lemma.
Lemma 7.4.
Let $f\in C^{1}(E)$ and consider $F\colon E^{r}_{X}\to\mathbb{R}$, $F(u_{1},\dots,u_{r}):=\sum_{i=1}^{r}f(u_{i})$. Then
$\nabla^{\mathrm{ver}}F(u)=\left(\nabla^{\mathrm{ver}}f(u_{1}),,\dots,\nabla^{\mathrm{ver}}f(u_{r})\right)\quad\forall u=(u_{1},\dots,u_{r})\in E^{r}_{X},$
where $\nabla^{\mathrm{ver}}F$ is considered with respect to $q_{r}$ and $\bar{g}$, while $\nabla^{\mathrm{ver}}f$ is considered with respect to $p$ and $g$.
Proof.
Fix $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$. For each $i\in\{1,2,\dots,r\}$ we put
$E_{i}:=\{0\}^{i-1}\times\mathrm{Ver}_{u_{i}}(p)\times\{0\}^{r-i}\subset T_{u}E^{r}_{X}.$
Taking orthogonal complements with respect to $\bar{g}|_{T_{u}E^{r}_{X}}$ and $g|_{T_{u_{i}}E}$, we observe that
$\left(\mathrm{Ver}_{u}(q_{r})\right)^{\perp}=\bigcap_{i=1}^{r}E_{i}^{\perp}.$
Apparently, $E_{i}^{\perp}=T_{u}E^{r}_{X}\cap(\prod_{j=1}^{i-1}T_{u_{j}}E\times(\mathrm{Ver}_{u_{i}}(p))^{\perp}\times\prod_{j=i+1}^{r}T_{u_{j}}E)$ for each $i\in\{1,2,\dots,r\}$, such that
$\left(\mathrm{Ver}_{u}(q_{r})\right)^{\perp}=T_{u}E^{r}_{X}\cap\prod_{i=1}^{r}(\mathrm{Ver}_{u_{i}}(p))^{\perp}.$
Let $\mathrm{pr}_{q_{r}}\colon T_{u}E^{r}_{X}\to T_{u}E^{r}_{X}$ and $\mathrm{pr}_{i}\colon T_{u_{i}}E\to T_{u_{i}}E$, $i\in\{1,2,\dots,r\}$, denote the orthogonal projections onto $\mathrm{Ver}_{u}(q_{r})$ and $\mathrm{Ver}_{u_{i}}(p)$, respectively, with respect to the chosen metrics. By definition and the above computation,
$\ker\mathrm{pr}_{q_{r}}=T_{u}E^{r}_{X}\cap\prod_{i=1}^{r}(\mathrm{Ver}_{u_{i}}(p))^{\perp}=T_{u}E^{r}_{X}\cap\prod_{i=1}^{r}\ker\mathrm{pr}_{i}=\ker\Big{(}\Big{(}\prod_{i=1}^{r}\mathrm{pr}_{i}\Big{)}\Big{|}_{TE^{r}_{X}}\Big{)}.$
We further observe that equation (6.2) can be rephrased by saying that the images of $\mathrm{pr}_{q_{r}}$ and $(\prod_{i=1}^{r}\mathrm{pr}_{i})|_{TE^{r}_{X}}$ coincide. Together with the equality of their kernels, this shows that
(7.1) $\mathrm{pr}_{q_{r}}=\Big{(}\prod_{i=1}^{r}\mathrm{pr}_{i}\Big{)}\Big{|}_{TE^{r}_{X}}.$
As orthogonal projections, the maps $\mathrm{pr}_{q_{r}}$ and $\mathrm{pr}_{i}$, $i\in\{1,2,\dots,r\}$, are self-adjoint with respect to the chosen metrics, so we obtain for all $v=(v_{1},\dots,v_{r})\in T_{u}E^{r}_{X}$ that
$\bar{g}\left(\nabla^{\mathrm{ver}}F(u),v\right)$ $=\bar{g}\left(\mathrm{pr}_{q_{r}}(\nabla^{\bar{g}}F(u)),v\right)=\bar{g}\left(\nabla^{\bar{g}}F(u),\mathrm{pr}_{q_{r}}(v)\right)$
[…]
$=\sum_{i=1}^{r}g(\nabla^{g}f(u_{i}),\mathrm{pr}_{i}(v_{i}))=\sum_{i=1}^{r}g(\mathrm{pr}_{i}(\nabla^{g}f(u_{i})),v_{i})$
$=\sum_{i=1}^{r}g(\nabla^{\mathrm{ver}}f(u_{i}),v_{i})=\bar{g}\left((\nabla^{\mathrm{ver}}f(u_{1}),\dots,\nabla^{\mathrm{ver}}f(u_{r})),v\right).$
Since $v$ was chosen arbitrarily, the claim immediately follows. ∎
Lemma 7.5.
Let $f\in C^{1}(E)$ and consider $F\colon E^{r}_{X}\to\mathbb{R}$, $F(u_{1},\dots,u_{r})=\sum_{i=1}^{r}f(u_{i})$. If $\nabla^{g}f\colon E\to TE$ is vertically proportional with respect to $p$ and $g$, then $\nabla^{\bar{g}}F\colon E^{r}_{X}\to TE^{r}_{X}$ is vertically proportional with respect to $q_{r}$ and $\bar{g}$.
Proof.
By definition, $F=f^{r}\circ i$, where $i\colon E^{r}_{X}\hookrightarrow E^{r}$ denotes the embedding, and one checks without difficulties that
$\nabla^{\bar{g}}F(u)=(Di_{u})^{*}(\nabla^{g^{r}}f^{r}(u))=(Di_{u})^{*}(\nabla^{g}f(u_{1}),\dots,\nabla^{g}f(u_{r}))$
for all $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$, where $(Di_{u})^{*}$ denotes the adjoint operator of $Di_{u}:T_{u}E^{r}_{X}\to T_{i(u)}E^{r}$ with respect to $\bar{g}$ and $g^{r}$. Thus, we compute for all $u=(u_{1},\dots,u_{r})$ that
$\|\nabla^{\bar{g}}F(u)\|_{\bar{g}}=\|(\nabla^{g}f(u_{1}),\dots,\nabla^{g}f(u_{r}))\|_{g^{r}}=\sqrt{\sum_{i=1}^{r}\|\nabla^{g}f(u_{i})\|_{g}^{2}},$
where the norms are induced by the metrics in their subscripts. Since $\nabla f$ is vertically proportional, there exists $C>0$ with $\|\nabla f(e)\|_{g}\leq C\cdot\|\nabla^{\mathrm{ver}}f(e)\|_{g}$ for all $e\in E$. Thus,
$\|\nabla F(u_{1},\dots,u_{r})\|_{\bar{g}}$ $\leq\sqrt{C^{2}\sum_{i=1}^{r}\|\nabla^{\mathrm{ver}}f(u_{i})\|_{g}^{2}}=C\cdot\|(\nabla^{\mathrm{ver}}f(u_{1}),\dots,\nabla^{\mathrm{ver}}f(u_{r}))\|_{g^{r}}$
$=C\cdot\|\nabla^{\mathrm{ver}}F(u_{1},\dots,u_{r})\|_{\bar{g}},$
where we used Lemma 7.4 for the final equality. ∎
If we apply Theorem 5.12 to a function as considered in the previous lemmas, we obtain the following result.
Theorem 7.6.
Let $f\in C^{1}(E)$ be bounded from below and consider
$F\colon E^{r}_{X}\to\mathbb{R},\qquad F(u_{1},\dots,u_{r})=f(u_{1})+\dots+f(u_{r}).$
Assume that $(E,g)$ is complete, that $(f,g)$ satisfies the PS condition, that $f(\operatorname{Crit}f)$ is discrete in $\mathbb{R}$ and that $\nabla^{g}f$ is vertically proportional with respect to $p$ and $g$. Then for all $\lambda\in\mathbb{R}$ it holds that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\max\left\{\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f)\mid\mu_{1}+\dots+\mu_{r}=\mu\right\}.$
Proof.
Since $(f,g)$ satisfies the PS condition, $(f^{r},g^{r})$ satisfies the PS condition on $E^{r}$. Since $E^{r}_{X}$ is a closed submanifold of $E^{r}$, it follows that $F=f^{r}|_{E^{r}_{X}}$ satisfies the PS condition as well. Moreover, $\nabla^{\bar{g}}F$ is vertically proportional with respect to $q_{r}$ by Lemma 7.4. Thus, it follows from Theorem 6.5.a) that
$\operatorname{\mathsf{TC}}_{r,p}(F^{\lambda})\leq\sum_{\mu\in(-\infty,\lambda]}\operatorname{\mathsf{TC}}_{r,p}(F_{\mu}\cap\operatorname{Crit}F).$
Since $\nabla^{g}f$ is vertically proportional, $f$ satisfies condition (A) and it follows from Proposition 7.3 that
$\operatorname{Crit}_{\mu}F$ $=F_{\mu}\cap\left(\operatorname{Crit}f\times_{X}\operatorname{Crit}f\times_{X}\dots\times_{X}\operatorname{Crit}f\right)$
$=\bigcup_{\mu_{1}+\mu_{2}+\dots+\mu_{r}=\mu}\operatorname{Crit}_{\mu_{1}}f\times_{X}\operatorname{Crit}_{\mu_{2}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f.$
Since $f$ is continuously differentiable, $\operatorname{Crit}_{\mu}f$ is closed in $E$ for each $\mu\in\mathbb{R}$, which yields that the latter is a union of closed subsets of $E^{r}_{X}$. One derives from the discreteness of $f(\operatorname{Crit}f)$ that the above is a finite union of pairwise disjoint closed subsets. Thus, by iteratively applying Proposition 5.9.d), we obtain that
$\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu}F)$ $=\operatorname{\mathsf{TC}}_{r,p}\Big{(}\bigcup_{\mu_{1}+\mu_{2}+\dots+\mu_{r}=\mu}\operatorname{Crit}_{\mu_{1}}(f)\times_{X}\operatorname{Crit}_{\mu_{2}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f\Big{)}$
$=\max\{\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{X}\operatorname{Crit}_{\mu_{2}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f)\mid\mu_{1}+\dots+\mu_{r}=\mu\}$
for all $\mu\in\mathbb{R}$. The claim immediately follows. ∎
We recall that in Theorem 4.2 we have established an upper bound on sequential TCs of sublevel sets by sequential subspace complexities that are not taken with respect to the whole domain of the function, but only with respect to the sublevel set itself. We want to establish an analogue of this result for sequential parametrized TCs and restrictions of fibrations of the form $p_{\lambda}:=p|_{f^{\lambda}}\colon f^{\lambda}\to X$. However, $p_{\lambda}$ will in general not be a fibration and will only have this property in very particular settings.
Remark 7.7.
The parametrized TC of maps which are not fibrations was studied in a general way by García Calcines in [GC23]. However, it remains an open question how much of the topological Lusternik-Schnirelmann theory that we used to establish our main estimates can be transferred to the more general setting.
We refer the reader to the next section for examples of situations in which this assumption is indeed satisfied.
Theorem 7.8.
Let $f\in C^{1}(E)$ be bounded from below and consider the function
$F\colon E^{r}_{X}\to\mathbb{R},\qquad F(u_{1},\dots,u_{r})=f(u_{1})+\dots+f(u_{r}).$
Assume that $(E,g)$ is complete, that $(f,g)$ satisfies the Palais-Smale condition and that $f(\operatorname{Crit}f)$ is discrete in $\mathbb{R}$. If $\lambda\in\mathbb{R}$ is chosen such that $f^{\lambda}$ is path connected, that
$p_{\lambda}:=p|_{f^{\lambda}}\colon f^{\lambda}\to X$
is a smooth fibration and that $\nabla^{g}f|_{f^{<\lambda+\delta}}$ is vertically proportional with respect to $p$ for some $\delta>0$, then
$\operatorname{\mathsf{TC}}_{r}[p_{\lambda}\colon f^{\lambda}\to X]$
$\leq\sum_{\mu\in(-\infty,r\lambda]}\max\left\{\operatorname{\mathsf{TC}}_{r,p_{\lambda}}(\operatorname{Crit}_{\mu_{1}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f)\mid\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda],\,\mu_{1}+\dots+\mu_{r}=\mu\right\}.$
Proof.
Let $Y_{f}\colon E\to TE$ and $Y_{F}\colon E^{r}_{X}\to TE^{r}_{X}$ be the pseudo-gradients of $f$ and $F$, respectively, that are constructed as in Proposition 6.4.a). Let $\varphi^{\mathrm{ver}}\colon E\times\mathbb{R}\to E$ and $\Phi^{\mathrm{ver}}\colon E^{r}_{X}\times\mathbb{R}\to E^{r}_{X}$ denote the flows of $Y_{f}^{\mathrm{ver}}$ and $Y_{F}^{\mathrm{ver}}$, respectively. One derives from Lemma 7.5 that
$Y_{F}^{\mathrm{ver}}(u_{1},\dots,u_{r})=(Y_{f}^{\mathrm{ver}}(u_{1}),\dots,Y_{f}^{\mathrm{ver}}(u_{r}))$
and thus
(7.2) $\Phi^{\mathrm{ver}}(u_{1},\dots,u_{r},t)=(\varphi^{\mathrm{ver}}(u_{1},t),\dots,\varphi^{\mathrm{ver}}(u_{r},t))\qquad\forall(u_{1},\dots,u_{r})\in E^{r}_{X},\,t\in\mathbb{R}.$
Since $f(\operatorname{Crit}f)$ is discrete, we can assume w.l.o.g. that $(\lambda,\lambda+\delta]$ does not contain a critical value of $f$. Put $U:=(f^{<\lambda+\delta})^{r}\cap E^{r}_{X}$. One derives from (7.2) that $\Phi^{\mathrm{ver}}(u,t)\in U$ for all $u\in U$ and $t\in[0,+\infty)$, such that $\Phi^{\mathrm{ver}}$ restricts to a flow
$\Psi:=\Phi^{\mathrm{ver}}|_{U\times[0,+\infty)}\colon U\times[0,+\infty)\to U.$
In other words, $Y_{F}^{\mathrm{ver}}|_{U}$ is forward complete. By Proposition 6.4.b), $(\varphi^{\mathrm{ver}}_{1},f)$ satisfies condition (D). Since $\partial U\subset(f_{\lambda+\delta})^{r}$ and since $\lambda+\delta$ is a regular value of $f$, it holds that $\partial U\cap\mathrm{Fix\;}\Phi=\emptyset$. Thus, by Lemma 4.1.b) and (7.2), $(\Psi_{1},F|_{U})$ satisfies condition (D).
Since $F(u)<r\lambda$ for each $u\in U$, it holds that $U=(F|_{U})^{r\lambda}$. Since $p_{\lambda}$ is a fibration, we can thus apply Theorem 6.5 to $\Psi_{1}$ and $F|_{U}$ and obtain that
$\operatorname{\mathsf{TC}}_{r}[p_{\lambda}\colon f^{\lambda}\to B]$ $=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}((f^{\lambda})^{r}_{X})=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}\left((F|_{U})^{r\lambda}\right)$
$\leq\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,p_{\lambda}}(F_{\mu}\cap\operatorname{Crit}F).$
From here on, the claim follows along the lines of the proof of Theorem 7.6. ∎
8. Equivariant fiber bundles and vertical proportionality
In contrast to the setting of sequential TCs in section 2 to 4, our Lusternik-Schnirelmann-type results for sequential parametrized setting in the last sections had the downside that some restrictive conditions have to be put on the chosen quantities, namely
  • the vertical proportionality of gradients in Theorem 6.5 and all of its consequences,
  • the fibration property of restrictions of fibrations to sublevel sets in Theorem 7.8.
In this section, we will study certain classes of examples of $G$-equivariant fiber bundles, where $G$ is a Lie group, in which these properties hold or can be derived from more tangible criteria.
Throughout this section we let $G$ be a connected finite-dimensional Lie group, $E$ and $B$ be smooth paracompact Banach manifolds and $p\colon E\to B$ be a smooth fiber bundle whose fibers are path-connected. We further assume that $E$ and $B$ are equipped with smooth left $G$-actions
$\Phi\colon G\times E\to E,\qquad\varphi\colon G\times B\to B,$
such that $p$ is $G$-equivariant with respect to these actions. Given a subset $A\subset B$ we further put $E|_{A}:=p^{-1}(A)$ and denote by $E_{b}:=E|_{\{b\}}=p^{-1}(\{b\})$ the fiber over $b$ for each $b\in B$.
Proposition 8.1.
Assume that the $G$-action $\varphi$ on $B$ is transitive and let $f\in C^{1}(E)$ be $G$-invariant.
  1. a)
    For each $a\in\mathbb{R}$, the restriction
    $p_{a}\colon f^{a}\to B,\qquad p_{a}:=p|_{f^{a}},$
    is a fiber bundle with typical fiber $E_{b}\cap f^{a}$ for some $b\in B$.
  2. b)
    Each $b\in B$ admits an open neighborhood $U\subset B$ and a local trivialization $\psi\colon E|_{U}\to U\times E_{b}$ of $p$, such that
    $(f\circ\psi^{-1})(b,x)=f(x)\qquad\forall b\in U,\ x\in E_{b}.$
Proof.
  1. a)
    Choose $b_{0}\in B$ and put $F:=E_{b_{0}}\cap f^{a}$. Since $\varphi$ acts transitively on $B$, the map
    $\pi_{\varphi}\colon G\to B,\qquad\pi_{\varphi}(g)=\varphi(g,b_{0}),$
    is a principal $G$-bundle, so there is an open neighborhood $U$ of $b_{0}$ which admits a local section $s\colon U\to G$ of $\pi_{\varphi}$. We consider the map
    $\psi\colon p_{a}^{-1}(U)\to U\times F,\qquad\psi(x):=(p(x),\Phi(s(p(x))^{-1},x)).$
    Note that $p_{a}^{-1}(U)=E|_{U}\cap f^{a}$. One checks that indeed $\Phi(s(p(x))^{-1},x)\in F$ for all $x\in p_{a}^{-1}(U)$. Moreover, the map $\psi$ is clearly continuous and one checks that a continuous inverse to $\psi$ is given by $\psi^{-1}(b,x)=\Phi(s(b),x)$, so that $\psi$ is indeed a local trivialization of $p_{a}$. Since $b$ was chosen arbitrarily, this shows that $p_{a}$ is a fiber bundle.
  2. b)
    Let $b_{0}$, $U$ and $\psi$ be given as in the proof of a). Then
    $f(\psi^{-1}(b,x))=f(\Phi(s(b),x))=f(x)\qquad\forall b\in U$
    by the $G$-invariance of $f$.
In Proposition 8.1.a) we have established a class of fiber bundles with the property that all of its restrictions to sublevel sets of a $G$-invariant function are again fiber bundles. To check in which of these settings we can apply Theorem 7.8, we next want to study the vertical proportionality of gradients of such functions.
Throughout the following, we put
$\Phi_{g}\colon E\to E,\quad\Phi_{g}(x):=\Phi(g,x)\quad\forall g\in G,$
$\Phi^{x}\colon G\to E,\quad\Phi^{x}(g):=\Phi(g,x)\quad\forall x\in E.$
We further
  • assume that $E$ admits and is equipped with a $\Phi$-invariant Riemannian metric, which we shall fiberwise denote by ,$_{x}$, $x\in E$,
  • assume that $B$ admits and is equipped with a Riemannian metric, for which $p\colon E\to B$ becomes a Riemannian submersion,
  • assume that $G$ admits and is equipped with a bi-invariant Riemannian metric.
[…]
Remark 8.2.
By a classical result of Milnor, see [Mil76, Lemma 7.5], a Lie group admits a bi-invariant metric if and only if it is isomorphic as a Lie group to a group of the form $K\times\mathbb{R}^{m}$, where $K$ is a compact Lie group and $m\in\mathbb{N}_{0}$. Thus, all Lie groups that we consider subsequently are of this form.
We want to establish a tangible criterion on a $C^{1}$-function on $f$ which implies the vertical proportionality of its gradient, for which we introduce some additional terminology.
Definition 8.3.
Let $f\in C^{1}(E)$ be bounded from below. We say that the $G$-action $\Phi$ is $f$-controlled if there exists a function $C\colon[\inf f,+\infty)\to(0,+\infty)$, such that
(8.1) $\|(D\Phi^{x})_{g}\|_{\mathrm{op}}\leq C(f(x))\quad\forall g\in G,\ x\in E.$
Remark 8.4.
As one easily sees, if $G$ is compact and if $f\in C^{1}(E)$ is such that $f^{\lambda}$ is compact for each $\lambda\in\mathbb{R}$, then $\Phi$ will be $f$-controlled. This holds in particular for any $f\in C^{1}(E)$ if both $G$ and $E$ are compact.
Lemma 8.5.
Let $f\in C^{1}(E)$ be $G$-invariant and bounded from below. Assume that $\varphi$ is a transitive $G$-action on $B$ and that $\Phi$ is $f$-controlled with respect to $C\colon[\inf f,+\infty)\to(0,+\infty)$. Let $\lambda\in\mathbb{R}$, let $U\subset B$ be open and let $b_{0}\in B$. Assume that there is a smooth local section $s\colon U\to G$ of $\pi_{\varphi}\colon G\to B$, $\pi_{\varphi}(g)=\varphi(g,b_{0})$. Then
$\|\nabla f(x)\|\leq(C(\lambda)\cdot\|Ds_{p(x)}\|_{\mathrm{op}}+1)\cdot\|\nabla^{\mathrm{ver}}f(x)\|\qquad\forall x\in f^{\lambda}\cap E|_{U}.$
Proof.
Since $f$ is $G$-invariant, $\nabla f$ is $G$-equivariant, i.e. it holds that
(8.2) $\nabla f(\Phi_{g}(x))=(D\Phi_{g})_{x}(\nabla f(x))\quad\forall g\in G,\ x\in E.$
Since the metric on $E$ is $G$-invariant and $p$ is $G$-equivariant, the differential of $\Phi_{g}$ maps vertical tangent spaces with respect to $p$ isomorphically onto one another for each $g\in G$. One derives from this observation and from (8.2) that
(8.3) $\nabla^{\mathrm{ver}}f(\Phi_{g}(x))=(D\Phi_{g})_{x}(\nabla^{\mathrm{ver}}f(x))\quad\forall g\in G,\ x\in E.$
Put $\eta\colon E|_{U}\to G$, $\eta(x):=(s(p(x)))^{-1}$, and $h\colon E|_{U}\to E$, $h(x):=\Phi(\eta(x),x)$. Since $f$ is $G$-invariant, it holds that $f(h(x))=f(\Phi(\eta(x),x))=f(x)$ for all $x\in E_{U}$, so we apply the chain rule to obtain that $Df_{x}=Df_{h(x)}\circ Dh_{x}$ for all $x\in E|_{U}$, or equivalently,
$\nabla f(x),v\right>=\left<\nabla f(h(x)),Dh_{x}(v)\right>\ \ \forall x\in E|_{U},\ v\in T_{x}E.$
As discussed in the proof of Proposition 8.1, it then holds that $h(E|_{U})\subset E_{b_{0}}$ and thus $Dh_{x}(v)\in T_{x}(E_{b_{0}})=\mathrm{Ver}_{x}(p)$ for all $x\in E|_{U}$ and $v\in T_{x}E$. We derive that
[…]
for all $x\in U$. Using the chain rule, we compute that
$Dh_{x}(v)$ $=D\Phi_{(\eta(x),x)}(D\eta_{x}(v),v)=(D\Phi^{x})_{\eta(x)}(D\eta_{x}(v))+(D\Phi_{\eta(x)})_{x}(v).$
for all $x\in E|_{U}$ and $v\in T_{x}E$. Inserting this into (8.4) yields
$\nabla f(x),v\right>_{x}$ =$(D\Phi_{\eta(x)})_{x}(\nabla^{\mathrm{ver}}f(x)),(D\Phi^{x})_{\eta(x)}(D\eta_{x}(v))+(D\Phi_{\eta(x)})_{x}(v)\right>_{h(x)}$
=$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(v)+v\right>_{x},$
where we have used the $G$-invariance of the metric. For $v=\nabla f(x)$ we obtain
$\|\nabla f(x)\|^{2}$ =$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\right>_{x}+\left<\nabla^{\mathrm{ver}}f(x),\nabla f(x)\right>_{x}$
=$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\right>_{x}+\|\nabla^{\mathrm{ver}}f(x)\|^{2},$
since $\nabla^{\mathrm{ver}}f(x),\nabla f(x)\right>=\left<\nabla^{\mathrm{ver}}f(x),\nabla^{\mathrm{ver}}f(x)\right>$, for each $x\in E|_{U}$. Using the Cauchy-Schwarz inequality and the $G$-invariance of the metric, we derive that
$\|\nabla f(x)\|^{2}$ $\leq\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}$
$\leq\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|(D\Phi^{x})_{\eta(x)}\|_{\mathrm{op}}\cdot\|D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
Since $\Phi$ is $f$-controlled with respect to $C$, we obtain
$\|\nabla f(x)\|^{2}$ $\leq C(f(x))\cdot\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}$
$\leq C(f(x))\cdot\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|D\eta_{x}\|_{\mathrm{op}}\cdot\|\nabla f(x)\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
If $\nabla f(x)\neq 0$, then dividing this inequality by $\|\nabla f(x)\|$ and using that $\|\nabla^{\mathrm{ver}}f(x)\|\leq\|\nabla f(x)\|$ yields
$\|\nabla f(x)\|\leq(C(\lambda)\cdot\|D\eta_{x}\|_{\mathrm{op}}+1)\|\nabla^{\mathrm{ver}}f(x)\|\quad\forall x\in f^{\lambda}\cap E|_{U}.$
Here, we use that the inequality is imminent if $\nabla f(x)=0$. It remains to consider $\|D\eta_{x}\|_{\mathrm{op}}$ for $x\in E|_{U}$. By definition, $\eta$ is given as $\eta=\mathrm{inv}\circ s\circ p$, where $\mathrm{inv}\colon G\to G$ denotes the inversion. Thus,
$\|D\eta_{x}\|_{\mathrm{op}}\leq\|D\mathrm{inv}_{s(p(x))}\|_{\mathrm{op}}\cdot\|Ds_{p(x)}\|_{\mathrm{op}}\cdot\|Dp_{x}\|_{\mathrm{op}}\qquad\forall x\in E|_{U}.$
Since $p$ is a Riemannian submersion, $\|Dp_{x}\|_{\mathrm{op}}=1$ for all $x\in E$. Let $\ell_{g},r_{g}:G\to G$ denote left- and right-multiplication, respectively, by $g\in G$. A simple computation shows that
$D\mathrm{inv}_{g}=-(Dr_{g^{-1}})_{e}\circ(D\ell_{g^{-1}})_{g}\qquad\forall g\in G.$
Since the chosen metric on $G$ is bi-invariant, this yields that $\|D\mathrm{inv}_{g}\|_{\mathrm{op}}=1$ for all $g\in G$. We derive that
$\|D\eta_{x}\|_{\mathrm{op}}\leq\|Ds_{p(x)}\|_{\mathrm{op}}\qquad\forall x\in f^{\lambda}\cap E|_{U}.$
Inserting this into the above inequality shows the claim. ∎
Proposition 8.6.
Assume that $B$ is a closed manifold and that $\varphi$ is a transitive $G$-action. Let $f\in C^{1}(E)$ be $G$-invariant and bounded from below. If $\Phi$ is $f$-controlled, then $\nabla f|_{f^{<\lambda}}$ is vertically proportional for each $\lambda\in\mathbb{R}$.
Proof.
Let $b_{0}\in B$ and consider again $\pi_{\varphi}\colon G\to B$, $\pi_{\varphi}(g)=\varphi(g,b_{0})$. Then $\pi_{\varphi}$ is a principal $G$-bundle, so every $b\in B$ has an open neighborhood $V_{b}$ which admits a smooth local section $\sigma_{b}\colon V_{b}\to G$ of $\pi_{\varphi}$. Since $B$ is a regular topological space, for each $b\in B$ there is an open neighborhood $U_{b}\subset B$ of $b$ with $\overline{U_{b}}\subset V_{b}$. By assumption $B$ is compact, so there exist finitely many $b_{1},\dots,b_{m}\in B$, for which $\{U_{b_{1}},\dots,U_{b_{m}}\}$ is an open cover of $B$. Put $U_{i}:=U_{b_{i}}$ and $s_{i}\colon U_{i}\to G$, $s_{i}:=\sigma_{b_{i}}|_{U_{i}}$ for each $i\in\{1,2,\dots,m\}$. Since each $\overline{U_{i}}$ is compact and $s_{i}$ has a smooth extension to $\overline{U_{i}}$, it follows that
$M_{i}:=\sup_{b\in U_{i}}\|(Ds_{i})_{b}\|_{\mathrm{op}}<+\infty.$
Let $C:[\inf f,+\infty)\to(0,+\infty)$ satisfy (8.1). Given $\lambda\in\mathbb{R}$, it then follows from Lemma 8.5 that
$\|\nabla f(x)\|\leq(C(\lambda)\cdot M_{i}+1)\cdot\|\nabla^{\mathrm{ver}}f(x)\|\quad\forall x\in f^{\lambda}\cap E|_{U_{i}},\ i\in\{1,2,\dots,m\}.$
Thus, with $C_{0}:=C(\lambda)\cdot\max\{M_{1},\dots,M_{m}\}+1$, it follows that $\|\nabla f(x)\|\leq C_{0}\|\nabla^{\mathrm{ver}}f(x)\|$ for all $x\in f^{\lambda}$, which shows that $\nabla f|_{f^{<\lambda}}$ is vertically proportional. ∎
This leads to the following Lusternik-Schnirelmann type result.
Theorem 8.7.
Let $r\in\mathbb{N}$ with $r\geq 2$. Assume that $B$ is a closed manifold and that $\varphi$ is a transitive $G$-action on $B$. Let $f\in C^{1}(E)$ be $G$-invariant and bounded from below. If $(E,\left<\cdot,\cdot\right>)$ is complete, if $(f,\left<\cdot,\cdot\right>)$ satisfies the PS condition and if $\Phi$ is $f$-controlled, then
$\operatorname{\mathsf{TC}}_{r}[p|_{f^{\lambda}}\colon f^{\lambda}\to B]$
$\leq\sum_{\mu\in(-\infty,r\lambda]}\max\left\{\operatorname{\mathsf{TC}}_{r,p_{\lambda}}(\operatorname{Crit}_{\mu_{1}}f\times_{B}\dots\times_{B}\operatorname{Crit}_{\mu_{r}}f)\mid\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda],\,\mu_{1}+\dots+\mu_{r}=\mu\right\}$
Proof.
This follows from Proposition 8.6 and Theorem 7.8. ∎
If both $E$ and $B$ are closed manifolds, the assumptions of Theorem 8.7 can be simplified drastically.
Corollary 8.8.
Let $r\in\mathbb{N}$ with $r\geq 2$. Assume that $E$ and $B$ are closed manifolds and that $\varphi$ is a transitive $G$-action on $B$. Then for any $G$-invariant $f\in C^{1}(E)$ it holds that
$\operatorname{\mathsf{TC}}_{r}[p\colon E\to B]\leq\sum_{\mu\in\mathbb{R}}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{B}\dots\times_{B}\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \mu_{1},\dots,\mu_{r}\in\mathbb{R},\,\sum_{i=1}^{n}\mu_{i}=\mu\Big{\}}.$
Proof.
Since $E$ is a closed manifold, $(E,\left<\cdot,\cdot\right>)$ is complete and $(f,\left<\cdot,\cdot\right>)$ satisfies the PS condition. Moreover, $\Phi$ is $f$-controlled, see Remark 8.4. Since $E$ is compact, $f$ attains its maximum $\lambda$, so that $p=p|_{f^{\lambda}}$. Then claim is thus an immediate consequence of Theorem 8.7. ∎
In the rest of this section, we want to apply Corollary 8.8 to unit tangent bundles of odd-dimensional spheres $\mathbb{S}^{2n-1}\subset\mathbb{R}^{2n}$. A general upper bound for the parametrized TCs of unit sphere bundles of vector bundles has been obtained by Farber and Weinberger in [FW23] and extended to sequential TCs by Farber and Paul in [FP24]. Using our approach via critical point theory, we will be able improve on these upper bounds for unit tangent bundles of odd-dimensional spheres and compute the exact values of their sequential parametrized TCs in certain cases.
In the following we fix $n\in\mathbb{N}$ and will use the canonical identification $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$ without always spelling it out explicitly. We further view $\mathbb{S}^{2n-1}$ as equipped with the metric induced by the standard metric on $\mathbb{R}^{2n}$ and denote its unit tangent bundle by
$p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}.$
Explicitly,
$U\mathbb{S}^{2n-1}=\{(x,v)\in\mathbb{S}^{2n-1}\times\mathbb{R}^{2n}\mid\langle x,v\rangle=0,\ \langle v,v\rangle=1\},$
[…]
$A\colon\mathbb{S}^{2n-1}\to U\mathbb{S}^{2n-1},\qquad A(x)=(x,ix),$
of $p$. We note that the scalar multiplication with $i$ is explicitly given by
(8.5) $ix=i(x_{1},\ldots,x_{2n})=(x_{2},-x_{1},x_{4},-x_{3},\ldots,x_{2n},-x_{2n-1})$
for $x=(x_{1},\ldots,x_{2n})\in\mathbb{R}^{2n}$. We can identify $U\mathbb{S}^{2n-1}$ with the Stiefel manifold of orthonormal $2$-frames in $\mathbb{R}^{2n}$,
$V_{2}(\mathbb{R}^{2n})=\{X\in M_{2n,2}(\mathbb{R})\mid X^{T}X=\mathbb{1}_{2}\}$
where $M_{k,m}(\mathbb{R})$ denotes the space of all real $k\times m$-matrices and where $\mathbb{1}_{2}\in M_{2,2}(\mathbb{R})$ denotes the identity matrix. Explicitly, we identify a unit tangent vector $(x,v)\in U\mathbb{S}^{2n-1}$ with the matrix $X=(x,v)$ where $x\in\mathbb{R}^{2n}$ and $v\in\mathbb{R}^{2n}$ are taken as the columns of $X$. The tangent space at an element $X\in U\mathbb{S}^{2n-1}$ is then identified with
$T_{X}U\mathbb{S}^{2n-1}=\{Y\in M_{2n,2}(\mathbb{R})\mid X^{T}Y+Y^{T}X=0\}.$
In particular, if we write $X\in U\mathbb{S}^{2n-1}$ as a matrix $X=(x_{1},x_{2})$ with columns $x_{1},x_{2}\in\mathbb{R}^{2n}$ and $Y\in T_{X}U\mathbb{S}^{2n-1}$ as $Y=(y_{1},y_{2})$ with columns $y_{1},y_{2}\in\mathbb{R}^{2n}$ we have
(8.6) $\langle x_{1},y_{1}\rangle=\langle x_{2},y_{2}\rangle=0\quad\text{as well as }\quad\langle x_{2},y_{1}\rangle+\langle x_{1},y_{2}\rangle=0.$
We consider the function
(8.7) $f\colon U\mathbb{S}^{2n-1}\to\mathbb{R},\quad f(X)=\langle A(x_{1}),x_{2}\rangle=\langle ix_{1},x_{2}\rangle\quad\text{for}\quad X=(x_{1},x_{2})\in U\mathbb{S}^{n}.$
Since $A$ is $SU(n)$-equivariant and since the Euclidean inner product is $SO(2n)$-invariant and thus $SU(n)$-invariant as well, $f$ is $SU(n)$-invariant. Both $U\mathbb{S}^{2n-1}$ and $\mathbb{S}^{2n-1}$ are closed manifolds, so we can apply Corollary 8.8 to this setting and obtain for each $r\geq 2$ that
(8.8) $\operatorname{\mathsf{TC}}_{r}[p]\leq\sum_{\mu\in\mathbb{R}}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{2n-1}}\dots\times_{\mathbb{S}^{2n-1}}\times\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \sum_{i=1}^{r}\mu_{i}=\mu\Big{\}},$
where $\operatorname{\mathsf{TC}}_{r}[p]:=\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}]$. We want to further compute this upper bound on $TC_{r}[p]$ and study the critical sets of $f$ in greater detail for this purpose.
Lemma 8.9.
Let $f\colon U\mathbb{S}^{2n-1}\to\mathbb{R}$ be given as in (8.7). Then $\operatorname{Crit}f=\operatorname{Crit}_{1}f\cup\operatorname{Crit}_{-1}f$, where
$\operatorname{Crit}_{1}f=\{(x,ix)\in U\mathbb{S}^{2n-1}\mid x\in\mathbb{S}^{2n-1}\}\quad\text{and}\quad\operatorname{Crit}_{-1}f=\{(x,-ix)\in U\mathbb{S}^{2n-1}\mid x\in\mathbb{S}^{2n-1}\}.$
Proof.
We compute the differential of $f$ at $X=(x_{1},x_{2})\in U\mathbb{S}^{2n-1}$ as
(8.9) $Df_{X}(Y)=\langle iy_{1},x_{2}\rangle+\langle ix_{1},y_{2}\rangle\qquad\forall Y=(y_{1},y_{2})\in T_{X}U\mathbb{S}^{2n-1}.$
Let $x\in\mathbb{S}^{2n-1}$. Evidently, $f(x,ix)=1$ and for each $Y=(y_{1},y_{2})\in T_{X}U\mathbb{S}^{2n-1}$ we have
$Df_{(x,ix)}(Y)=\langle ix,iy_{1}\rangle+\langle ix,y_{2}\rangle=-\langle x,y_{1}\rangle+\langle ix,y_{2}\rangle=0$
by equation (8.6). Hence, $(x,ix)\in\operatorname{Crit}_{1}f$ and similarly one shows that for each $x\in\mathbb{S}^{2n-1}$ it holds that $(x,-ix)\in\operatorname{Crit}_{-1}f$. Thus, we have shown that
$\{(x,ix)\in U\mathbb{S}^{2n-1}\,|\,x\in\mathbb{S}^{2n-1}\}\cup\{(x,-ix)\in U\mathbb{S}^{2n-1}\,|\,x\in\mathbb{S}^{2n-1}\}\subset\operatorname{Crit}f.$
To show the converse inclusion, let $X=(x_{1},x_{2})\in U\mathbb{S}^{2n-1}$ with $x_{2}\notin\{ix_{1},-ix_{1}\}$. Set $Y=(0,w)$ with $w=ix_{1}-\langle ix_{1},x_{2}\rangle x_{2}$. One verifies that indeed $Y\in T_{X}U\mathbb{S}^{2n-1}$ and that
$Df_{X}(Y)=1-\langle ix_{1},x_{2}\rangle^{2}\neq 0.$
Hence, $X\not\in\operatorname{Crit}f$, which shows that any critical point of $f$ is of the form derived above. ∎
Let $r\in\mathbb{N}$ with $r\geq 2$. We have seen in the previous Lemma that the critical values of $f$ are $1$ and $-1$, so the fiber products of the critical manifolds occurring in (8.8) are of the form
$\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{2n-1}}\operatorname{Crit}_{\mu_{2}}f\times_{\mathbb{S}^{2n-1}}\ldots\times_{\mathbb{S}^{2n-1}}\operatorname{Crit}_{\mu_{r}}f\quad\text{with}\quad\mu_{j}\in\{+1,-1\},\,\,j\in\{1,2,\ldots,r\}.$
The sums of critical values that occur as values of $\mu$ on the right-hand side of (8.8) must further lie in $\{-r,-r+2,\dots,r-2,r\}$. Thus, it follows from (8.8) that
(8.10) $\operatorname{\mathsf{TC}}_{r}[p]\leq\sum_{k=0}^{r}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{2n-1}}\dots\times_{\mathbb{S}^{2n-1}}\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \sum_{j=1}^{r}\mu_{j}=2k-r\Big{\}}.$
We want to explicitly compute the sequential parametrized subspace complexities that occur on the right-hand side of this inequality in the case of even $n$, i.e. for spheres of dimension $14m-$ 1 for some $m\in\mathbb{N}$. Eventually, we will show the following.
Theorem 8.10.
For each $m\in\mathbb{N}$, it holds that
$\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}]=r+1.$
For this purpose, we will use the canonical identification $\mathbb{R}^{4}\cong\mathbb{H}$ with $\mathbb{H}$ denoting the quaternions, and consider $\mathbb{R}^{4m}\cong\mathbb{H}^{m}$ as a left $\mathbb{H}$-module. The quaternionic units $i,j,k\in\mathbb{H}$ can be chosen in such a way that multiplication with $i$ is again given by equation (8.5) and that left-multiplication with $j\in\mathbb{H}$ on $\mathbb{R}^{4m}$ is given by
$j(x_{1},\ldots,x_{4m})=(-x_{4},-x_{3},x_{2},x_{1},-x_{8},-x_{7},x_{6},x_{5},\ldots,-x_{4m},-x_{4m-1},x_{4m-2},x_{4m-3}).$
In particular, it holds that
(8.11) $\langle x,jx\rangle=\langle ix,jx\rangle=0\quad\text{and}\quad\|jx\|^{2}=\|x\|^{2}\qquad\forall x\in\mathbb{R}^{4m}.$
Lemma 8.11.
Let $m\in\mathbb{N}$ and let $f\colon\mathbb{S}^{4m-1}\to\mathbb{R}$ be given as in (8.7). Let $\mu_{1},\ldots,\mu_{r}\in\{+1,-1\}$ and consider the manifold
$\mathcal{C}=\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{2}}f\times_{\mathbb{S}^{4m-1}}\ldots\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{r}}f.$
Then $\operatorname{\mathsf{TC}}_{r,p}(\mathcal{C})=1$.
Proof.
Consider the map $B\colon\mathbb{S}^{4m-1}\to U\mathbb{S}^{4m-1}$, $B(x)=jx$. We obtain from (8.11) that $B$ is well-defined and satisfies $\langle A(x),B(x)\rangle=0$ for all $x\in\mathbb{S}^{4m-1}$.
Let $u=((x,v_{1}),\ldots,(x,v_{r}))\in\mathcal{C}$, so that $v_{j}\in\{ix,-ix\}$ for all $j\in\{1,2,\ldots,r\}$. We define a path $\sigma_{u}\colon I\to U\mathbb{S}^{n}$ as follows. For all $j\in\{0,1,\ldots,r-2\}$ and $t\in[\frac{j}{r-1},\frac{j+1}{r-1}]$ we put
$\sigma_{u}(t)=\begin{cases}(x,v_{j+1})&\text{if }v_{j+2}=v_{j+1},\\ \big{(}x,\cos(\pi((r-1)t-j))v_{j+1}+\sin(\pi((r-1)t-j))B(x)\big{)}&\text{if }v_{j+2}=-v_{j+1}.\end{cases}$
Clearly, $\sigma_{u}$ is a continuous path whose image lies in $p^{-1}(\{x\})$. One further checks that the fibration $\Pi_{r}\colon(U\mathbb{S}^{4m-1})^{I}_{\mathbb{S}^{4m-1}}\to(U\mathbb{S}^{4m-1})^{r}_{\mathbb{S}^{4m-1}}$ satisfies
$\Pi_{r}(\sigma_{u})=((x,v_{1}),(x,v_{2}),\ldots,(x,v_{r})).$
Moreover, the map $s\colon\mathcal{C}\to(U\mathbb{S}^{4m-1})_{\mathbb{S}^{4m-1}}^{I}$, $s(u)=\sigma_{u}$, is continuous and it follows that $s$ is a continuous sequential parametrized motion planner on $\mathcal{C}$. Since $\mathcal{C}$ is closed in $(U\mathbb{S}^{4m-1})^{r}_{\mathbb{S}^{4m-1}}$ and since $U\mathbb{S}^{4m-1}$ and $\mathbb{S}^{4m-1}$ are closed connected manifolds, hence path-connected ANRs, it follows from Lemma 5.8 that $s$ extends to a continuous sequential parametrized motion planner on an open neighborhood of $\mathcal{C}$. Hence, $\operatorname{\mathsf{TC}}_{r,p}(\mathcal{C})=1$. ∎
Proof of Theorem 8.10.
It follows from Lemma 8.11 that each summand of the form
$\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{4m-1}}\dots\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \sum_{j=1}^{r}\mu_{j}=2k-r\Big{\}}$
occurring on the right-hand side of (8.10) equals one. Since there are precisely $r+1$ of those summands, this yields
$\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}]\leq r+1.$
It is shown in [FP22, p. 9] that the $r$-th sequential parametrized TC of a fibration is bounded from below by the $r$-th sequential TC of its fiber. Each fiber of $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ is homeomorphic to $\mathbb{S}^{4m-2}$, so we derive that
$\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}]\geq\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{4m-2})=r+1,$
using the computation of sequential TCs of spheres from [Rud10, Section 4]. Combining the lower and the upper bound shows the claim. ∎
Remark 8.12.
We note that the unit tangent bundle $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ is fiber homotopy equivalent to a trivial fibration if and only if $m\in\{1,2\}$, so that Theorem 8.10 can not be derived by elementary arguments for $m\geq 3$.
More precisely, as shown in [JW54, Theorem 1.1], $U\mathbb{S}^{4m-1}$ has the homotopy type of $\mathbb{S}^{4m-1}\times\mathbb{S}^{4m-2}$ if and only if there exists an element of Hopf invariant one in $\pi_{8m-1}(\mathbb{S}^{4m})$. By [Ada60] this is the case only for $m\in\{1,2\}$. Thus, $p$ is not fiber homotopy trivial for $m\geq 3$.